### v2

# AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library

Abstract
Optimization modeling enables critical decisions across industries but remains hard to automate: informal language must be mapped to precise mathematical formulations and executable solver code, while prior LLM approaches either rely on brittle prompting or costly retraining with limited generalization. We present AlphaOPT, a self-improving experience library that enables an LLM to learn from limited demonstrations (i.e, even answers alone without gold-standard program) and solver feedback without annotated reasoning traces or parameter updates. AlphaOPT operates a continual two-phase cycle: (i) a Library Learning phase that reflects on failed attempts, extracts solver-verified, structured insights as $\{\textit{taxonomy},\ \textit{condition},\ \textit{explanation},\ \textit{example}\}$; and (ii) a Library Evolution phase that diagnoses retrieval misalignments and refines the applicability conditions of stored insights, improving transfer across tasks. This design (1) learns efficiently from limited demonstrations without curated rationales, (2) expands continually without costly retraining by updating the library rather than model weights, and (3) makes knowledge explicit and interpretable for human inspection and intervention. Experiments show that AlphaOPT steadily improves with more data (65% $\rightarrow$ 72% from 100 to 300 training items) and surpasses the strongest baseline by 7.7% on the out-of-distribution OptiBench dataset when trained only on answers. AlphaOPT code and data are available at https://github.com/Minw913/AlphaOPT.
1 INTRODUCTION
Optimization models support critical decision-making in finance, manufacturing, marketing, transportation, and logistics (AhmadiTeshnizi et al., 2023; Bertsimas & Tsitsiklis, 1997; Ramamonjison et al., 2022). Beyond improving efficiency, automating the optimization workflow lowers the barrier to operations research expertise in industry, enabling non-experts to prototype faster, iterate on formulations, and deploy solver-backed decisions at scale. Yet this process has long been challenging, as informal and often ambiguous specifications must be mapped to precise, domain-specific formulations and paired with appropriate code and solvers, creating major bottlenecks for end-to-end automation (Jiang et al., 2025).
Advances in large language models (LLMs) make this vision increasingly feasible: they can parse natural language requirements (Ouyang et al., 2022), generate executable programs (Nijkamp et al., ; Jimenez et al., 2024), and orchestrate downstream tools (Qin et al., 2024). Two main lines of work have emerged. Prompt-based systems steer general LLMs with structured prompts and tool use (Xiao et al., 2023; Thind et al., 2025; AhmadiTeshnizi et al., 2024; Zhang & Luo, 2025). Fine-tuning approaches adapt models on domain corpora and benchmarks (Huang et al., 2025; Yang et al., 2024). Despite this progress, both families face limitations: prompt-based systems stop improving once they run out of fixed templates, and they are fragile to small wording changes and shifts in the domain; fine-tuned models require costly retraining and, critically, most benchmarks and datasets in the community (e.g., NLP4LP (AhmadiTeshnizi et al., 2024), MAMO (Huang et al., 2024), IndustryOR (Huang et al., 2025)) contain only programs/solutions rather than the intermediate reasoning that governs modeling choices, thereby limiting the generalizability of fine-tuning approaches. This motivates a new learning paradigm for optimization formulation: instead of relying solely on prompts or retraining, LLMs should continually improve by accumulating, refining, and reusing solver-verified modeling insights.
Figure 1: AlphaOPT identifies and retrieves relevant insights to guide problem solving. In this example, it avoids the common mistake of minimizing the sum of process times and instead introduces an auxiliary variable to correctly minimize the makespan, leading to the correct solution.
We propose AlphaOPT, a self-improving framework that builds and refines a structured library of solver-verified insights for optimization formulation, as exemplified in Figure 1. Each insight encodes a reusable modeling rule in the form of a 4-tuple (taxonomy, (applicability) condition, explanation, example), which specifies not only what to reuse but also when and why it applies. We remark that our library learning framework does not require backpropagation to update framework parameters and can be regarded as the evolutionary mechanism. AlphaOPT improves through a continual two-phase cycle. Library Learning acquires new insights from both gold programs (when available) and solver-verified answer-only supervision, organizing them into a dynamically updating hierarchical taxonomy. Library Evolution then diagnoses misalignments between tasks and insight applicability, and refines conditions using aggregate evidence, ensuring that insights remain neither too narrow nor overly general. This design yields a principled optimization view: library construction corresponds to maximizing expected task success induced by task–insight matching while regularizing size to maintain efficiency and prevent redundancy.
We conduct quantitative experiments across multiple benchmarks and baselines, as well as qualitative analyses of the learned library. The results show that, compared to prompt-based or fine-tuning approaches, AlphaOPT (1) learns efficiently from limited demonstrations (i.e., it can learn from answers without recalling formulation) without requiring annotated reasoning traces or even gold-standard programs, (2) achieves stronger out-of-distribution generalizability and more consistent continual growth than learning-based methods, and (3) makes knowledge explicit and interpretable for human inspection and involvement.
Beyond these advantages, AlphaOPT also achieves state-of-the-art performance on multiple benchmarks. These results demonstrate the efficacy and potential of self-improving experience-library learning for optimization formulation, paving the way toward more challenging settings, such as efficient program formulation and large-scale optimization.
Our main contributions are threefold:
  • A library learning framework that learns from answers only. We propose the first experience-library learning framework for natural language optimization formulation tasks, formally grounded in a mathematical view. The system can learn solely from answers, without requiring gold-standard programs.
  • A reusable and interpretable experience library. We construct the first solver-verified library of structured modeling insights for LLM systems, designed to be reusable across tasks and explicitly interpretable for reliable transfer in operations research domains.
  • State-of-the-art out-of-distribution generalization. AlphaOPT achieves strong generalization beyond training distributions, attaining state-of-the-art performance on LogiOR and OptiBench benchmarks.
2 Related Work
LLMs for Solving Optimization Problems. Related work can be categorized into prompt-based and learning-based approaches. Prompt-based methods guide reasoning and modeling through multi-step prompts using proprietary LLMs (AhmadiTeshnizi et al., 2024; Xiao et al., 2023). AhmadiTeshnizi et al. (2023) first introduced OptiMUS, demonstrating how LLMs can generate optimization formulations from natural language descriptions, and OptiMUS-0.3 (AhmadiTeshnizi et al., 2024) extends this line of work to large-scale instances, introducing retrieval-augmented prompting and solver-integrated verification to improve scalability.
In contrast, learning-based methods construct training datasets and apply instruction tuning to open-source LLMs. Huang et al. (2025) proposed a semi-automated pipeline to synthesize training data, which is then used to fine-tune an open-source ORLM model. LLMOPT (Jiang et al., 2025) combines both paradigms by modeling optimization with five elements and fine-tuning on expert-annotated data via multi-instruction SFT. More recently, ORThought (Yang et al., 2025a) introduced the LogiOR benchmark and an expert-guided chain-of-thought framework, providing a systematic dataset and evaluation pipeline for optimization tasks that require harder logic. In terms of multi-agent design, Xiao et al. (2023) employs a collaborative multi-expert framework to enhance reasoning, Zhang & Luo (2025) integrates sandbox-based code execution and self-repair/self-verification.
Several benchmarks now exist that cover LP, MILP, NLP, and other optimization problem types (Xiao et al., 2023; AhmadiTeshnizi et al., 2024; Huang et al., 2025; Yang et al., 2024). Yet, none of the prior work has investigated strengthening LLMs’ optimization capabilities by learning and reusing structured modeling experience.
Decision-making tasks with Library Learning. Library Learning refers to the process where reusable patterns or modules are automatically extracted from past experiences to improve future problem-solving. These experiences include concrete trajectories or demonstrations, as well as abstracted rules generalized from successful or failed attempts (Zhao et al., 2024; Mu et al., 2025; Feng et al., 2025; Wang et al., ; Zhu et al., 2023). In terms of experience improvement, Zhao et al. (2024) and Mu et al. (2025) leverage an LLM to prune the library by checking if a newly added insight duplicates or conflicts with existing insights, or merges and generalizes from those overlapping insights. Feng et al. (2025) uses check functions to ensure that LLM-translated action sequences remain within the generalization boundaries of the original experiences.
LLM-driven Evolutionary Methods. Recent LLM-driven evolutionary frameworks have achieved remarkable advances in scientific discovery, showcasing LLM’s capacity for broad generative exploration on solutions. Romera-Paredes et al. (2024) introduces FunSearch, a genetic programming driven by LLMs to search for feasible or optimal solutions of mathematical problems. AlphaEvolve (Novikov et al., 2025) extends the FunSearch system to provide the ability to perform multiobjective optimization using rich forms of natural-language context and feedback. Grayeli et al. (2024) applies LLMs to discover abstract concepts from high-performing hypotheses, combining symbolic regression with LLM-guided exploration within a concept library. ReEvo (Ye et al., 2024) frames LLMs as hyper-heuristics with a reflective evolution mechanism, enabling the generation and iterative refinement of heuristics across multiple NP-hard problems. HeurAgenix (Yang et al., 2025b) further develops this direction by evolving a pool of heuristics and dynamically selecting the most suitable one for each problem state, highlighting LLMs’ role in adaptive heuristic discovery. Besides, LLM-guided evolution has also found use in discovering heuristics for combinatorial optimization (Liu et al., ).
3 Methodology
Optimization tasks arrive with diverse natural-language descriptions, yet they share recurring modeling rules that activate under identifiable conditions. We identify three major challenges in building reliable systems that leverage LLMs to formulate and solve optimization problems using existing technologies and resources. First, gold-standard programs are scarce and may contain annotation errors (Jiang et al., 2025; Yang et al., 2025a), while datasets with only answer labels remain underutilized (Huang et al., 2024; 2025; Lu et al., 2025). Second, fine-tuned models (Huang et al., 2025; Jiang et al., 2025) struggle to generalize because the crucial when-to-apply-what knowledge is weakly represented (or missing) in training data; they can mimic syntax without mastering applicability. Third, the performance of prompt-based agent systems AhmadiTeshnizi et al. (2023); Xiao et al. (2023); Yang et al. (2025a) stagnates as the number of exemplars increases: they rely on human empirical curation and lack the capacity to adapt or to continually learn from larger datasets.
We propose AlphaOPT, an experience-library learning framework that learns from both gold programs (when available) and answer-only supervision. AlphaOPT iteratively builds a structured, solver-verified repository of reusable insights with explicit applicability conditions and evolves these conditions at the population level to improve generalization while avoiding redundancy. This two-stage design is described in Section 3.1. In Section 3.2, we provide a mathematical interpretation that frames library construction as maximizing task success with a size regularizer. In Section A, we compare our method with prior works on learning from experience and self-evolving problem-solving agents.
3.1 AlphaOPT Framework
Figure 2: AlphaOPT builds and refines an experience library over multiple iterations. Each iteration consists of a library learning stage, which produces insights from attempts to solve tasks, and a library refinement stage, which adjusts the applicability of insights to avoid being too narrow or too general.
The framework incrementally learns a structured library of experiences over iterations until a stopping criterion indicates that the current model can no longer make meaningful improvements. As illustrated in Figure 2, each iteration consists of two complementary phases that form a continual cycle of acquisition and refinement. The first phase, Library Learning, extracts insights from individual tasks under either gold-program or answer-only supervision while minimizing redundancy. The second phase, Library Evolution, diagnoses misalignments between insights and tasks and refines applicability conditions to enhance generalization while reducing confusion caused by overgeneralization. The design follows three guiding principles: it is failure-driven (every error becomes a learning opportunity), locally validated (an insight must solve its source task before being admitted), and compact yet generalizable (redundant insights are merged and conditions refined to prevent uncontrolled growth that hinders retrieval and execution). The prompts for all LLM modules are provided in Appendix E.
Figure 3: The overall workflow of library learning and evolution. The left panel depicts two complementary flows: library learning, which extracts new insights by generating and consolidating insights from failed optimization attempts, and library diagnosis, which analyzes interactions between failed tasks and retrieved insights to collect negative and unretrieved cases for refinement. The right panel illustrates library refinement, where the LLM refines each insight’s applicability conditions, validates them via retrieval replay, and integrates the updated insights back into the library.
3.1.1 Library Learning
The objective of this stage is to generate reusable insights as structured 4-tuples (Taxonomy, Condition, Explanation, Example) and organize them in a hierarchical taxonomy for efficient retrieval, while minimizing redundancy in the library. The left panel of Figure 3 (see the Library learning flow label) illustrates the workflow for this stage.
Insight Extraction, Representation, and Supervision Mode.
Insights can be learned from either problems with a gold-standard program or with the answer alone. For each task, the system first constructs a mathematical formulation, then generates an executable solver program, and invokes the solver. When the library is non-empty, both steps are guided by retrieved insights. If the generated program does not achieve the correct optimal value, two supervision modes are used to guide the generation of new insights. When a gold program is available, the system compares the candidate formulation and program against the reference, diagnosing discrepancies (e.g., missing variables, misformulated constraints, incorrect objective terms) and distilling them into insights. When only the answer (i.e., final optimal objective) is provided, the system performs solver-guided self-exploration: it iteratively proposes executable programs, reuses prior failures as context, and receives verification from the solver. Once a program achieves its correct objective, it is treated as a proxy for the gold standard in anchor insight extraction. Before being stored in the library, each insight is locally verified by reapplying it to its source task to ensure that it resolves the original failure. In addition, to mitigate stochastic successes that could obscure useful lessons, we conduct three independent trials per task, allowing errors from probabilistic generation to serve as signals for learning.
Each insight is represented as a structured 4-tuple: Taxonomy, hierarchical labels for indexing and retrieval; Condition, an explicit description of the applicability signals in the problem; Explanation, the underlying principle of applying this insight; and Example, a concrete demonstration such as a mathematical constraint or code snippet.
Library Storage and Retrieval.
Insights are stored in a dynamically updated hierarchical taxonomy organized into three main tracks: Domain Modeling (problem-specific structures and assumptions), General Formulation (reusable mathematical patterns), and Code Implementation (solver-specific coding practices). Under each track, insights are further classified with two-level labels, where Level-1 captures a broad category and Level-2 refines it into a more specific subcategory. The taxonomy is initialized with few-shot labels and expands online: each new insight is either mapped to an existing category or, if no suitable label exists, prompts the LLM to propose new Level-1 or Level-2 labels. Each label is also assigned a condition, written by the LLM, that specifies when the category should be retrieved. When storing insights, to reduce redundancy, the LLM also checks whether a similar insight already exists and performs merging when appropriate. To align a target task with relevant insights, we employ a two-step LLM-driven retrieval procedure: Quick label matching, then full applicability check. The system first scans the taxonomy dictionary to identify labels that are potentially relevant to the context of the tasks. For example, Level-2 label such as Fixed Charge (Big-M Linking) will be probably detected when the problem description specifies that service or flow is allowed only if a facility is opened. After candidate labels are identified, the system rigorously evaluates each associated insight by examining its condition, and only the most applicable insights are retained.
During solution generation, retrieved insights from the Domain Modeling and General Formulation tracks guide the construction of the mathematical model, while insights from the Code Implementation track guide solver-code generation. This two-step procedure ensures that insights are applied appropriately, while the hierarchical taxonomy provides an extensible structure for organizing and retrieving insights as the library grows.
Operational Flow.
Training proceeds in an online regime over minibatches of data, starting from an empty library. For each batch, the system retrieves candidate insights, generates and executes programs, and upon failures extracts insights and immediately commits those that pass local self-verification, allowing newly added insights to benefit subsequent tasks and preventing the generation of repetitive insights. To reduce redundancy, tasks are clustered and reordered by problem type and semantic similarity, and overlapping insights within a batch are merged prior to integration. The process iterates until overall accuracy plateaus, at which point the library is archived and used for evaluation.
In implementation, for the sake of efficiency, training follows two coordinated data flows. The first processes minibatches of tasks in parallel for insight extraction. The second maintains a centralized queue of all generated insights, storing them into the library sequentially. This queue does not allow asynchronous updates, as concurrent modifications could lead to conflicts if two insights attempt to update the library simultaneously. This design balances parallelism in problem-solving with strict serialization in library updates, ensuring both efficiency and consistency.
3.1.2 Library Evolution
Figure 4: A locally verified initial condition (light-grey dashed circle) is refined into a broader applicability boundary (solid black) through evolutionary prompt optimization guided by the aggregate metric.
While Library Learning expands the repository of insights, Library Evolution aims to transform task-specific lessons into broadly applicable knowledge. Since each insight’s applicability is defined by a condition induced from a specific task, early conditions are often too narrow (failing to trigger on relevant tasks) or too broad (causing misretrieval). Left unchecked, these misalignments lead to missed opportunities or systematic failures. Library Evolution counters this with a diagnostic–refinement cycle: it detects misaligned insights, aggregates evidence across tasks, and refines conditions at the end of each iteration. The refinement is guided by an aggregate metric rather than ad-hox fixes. As illustrated in Figure 4, library refinement can be understood as adjusting each insight’s condition toward the correct retrieval boundary in the problem space.
Diagnostic: Library Diagnosis.
After each training round, we trace failed tasks and analyze their interaction with the library. The diagnostic agent partitions the relationship between each insight $i$ and its associated tasks into three disjoint categories: $\Pi(i)=\{\text{Positive}:S_{i}^{+},\text{Negative}:S_{i}^{-},\text{Unretrieved}:S_{i}^{u}\}$ where $S_{i}^{+}$ contains tasks where the insight was applicable and contributed to the correct formulation, $S_{i}^{-}$ contains tasks where it was misleading and degraded performance, and $S_{i}^{u}$ contains tasks where it was not retrieved but would have been beneficial. By maintaining these partitions across iterations, the system continuously builds a performance profile for each insight. If a failed task is subsequently solved after removing a misleading (negative) insight or by injecting a previously unretrieved one, the system attributes the failure to condition misalignment rather than lack of knowledge, thus avoiding redundant insight generation. Unretrieved tasks are identified by first comparing the model’s generated program with the ground-truth (or a reference program from self-exploration) to locate discrepancies. These discrepancies guide the search for candidate insights, which are then verified for their ability to resolve the errors. Verified insights are flagged as relevant but unretrieved, allowing the system to diagnose retrieval gaps without resorting to intractable combinatorial search. The left panel of Figure 3 (see the Library diagnosis flow label) illustrates the workflow for this stage.
Evolver: Library Refinement.
Building on the diagnosis, the Evolver agent refines insights in two steps: condition refinement and refinement verification. First, for each diagnosed insight, the agent strengthens or prunes its applicability condition. Negative tasks contribute explicit inapplicability clauses (e.g., constraints or contexts that block use), while unretrieved tasks highlight missing applicability signals. The Evolver then proposes multiple refinement strategies (e.g., adding preconditions, introducing keyword anchors, merging overlapping triggers) and produces candidate conditions with the goal of preserving correct cases, eliminating mismatches, and recovering previously missed tasks. Then, each candidate condition replaces the original and is tested over the union $R_{i}=S_{i}^{+}\cup S_{i}^{-}\cup S_{i}^{u}$. A performance score
$p_{i}=\frac{|\text{kept positives}|+|\text{corrected negatives}|+|\text{recovered unretrieved}|}{|R_{i}|}$
quantifies improvement. Here, “kept positives” are tasks that still remain correctly retrieved after refinement; “corrected negatives” are tasks that were misled by the insight before and no longer retrieved; and “recovered unretrieved” are tasks that become correctly retrieved after refinement. We accept refinements that increase $p_{i}$ and keep the one with the highest $p_{i}$. The workflow for this stage is illustrated in the right panel of Figure 3.
3.2 Optimization Perspective
The framework can be viewed as an iterative solution to the optimization problem in the library space. Let $\mathcal{L}$ denote a candidate library and $\mathcal{T}$ the distribution of the optimization problems we want to solve. The objective is to maximize task success while penalizing library complexity to mitigate retrieval inefficiency and long-context degradation in LLM inference:
$\max_{\ell\in\mathcal{L}}\;\;\mathbb{E}_{t\sim\mathcal{T}}\!\left[\mathrm{Success}(t\mid\ell)\right]\;-\;\lambda\,\Omega(\ell).$
where $\mathrm{Success}(t\mid\ell)$ indicates whether $\ell$ enables the system to produce a program that achieves the correct optimal objective for task $t$, and $\Omega(\ell)$ quantifies library complexity (e.g., number of insights or redundancy-adjusted size). Under our problem design—bounded and continuous property of $\text{Success}(\cdot)$ and $\Omega(\cdot)$, sufficient exploration under solver verification, and bounded merging—the refinement dynamics converge to a locally optimal library. In Appendix C, we provide a conceptual sketch showing that convergence holds: As refinement in the second phase strictly improves the first term, while verified merging in the first phase reduces the second term without diminishing the first, sufficient exploration combined with iterative cycles of library learning and evolution ensures convergence to a local optimum. Given the inherent ambiguity of natural language and stochasticity in LLM outputs, we present this perspective not as a strict theorem but as a principled justification for the acquisition–refinement design and the redundancy-reduction operations.
4 EXPERIMENTS
Our experiments are designed to reflect the requirements that arise in real-world optimization and operations research (OR) applications. In these settings, methods are expected not only to perform well on standard benchmarks, but also to transfer across domains, to remain effective when limited supervision is available, to improve steadily as more data becomes available, and to offer results that can be inspected and audited. We therefore organize our evaluation around four questions: (1) How well does the method generalize across domains? (2) Can it learn effectively with limited supervision? (3) Does performance improve consistently with more training data? (4) How does it compare overall with strong baselines? Finally, we examine the interpretability of the insight library to assess whether the outputs are understandable and actionable to practitioners.
4.1 Experimental Setup
Our experiments are conducted on a dataset of 454 problem instances, aggregated from four real-world optimization and operation task datasets, namely the NLP4LP (AhmadiTeshnizi et al., 2024), NL4OPT (Ramamonjison et al., 2022), IndustryOR (Huang et al., 2025), MAMO (ComplexLP) (Huang et al., 2024), with any invalid entries discarded. These collections span various formulation types and originate from diverse sources, including academic papers, textbooks, and real‐world industry scenarios. Detailed descriptions of these datasets are provided in Appendix B.
We perform stratified sampling within each dataset, randomly partitioning $70\%$ for training and $30\%$ for testing. We maintain a strict separation between training and test data. The experience library is constructed only from training tasks, ensuring that training-derived insight examples do not leak into the test set. To assess out-of-distribution generalization, we additionally evaluate on LogiOR (Yang et al., 2025a) and Optibench (Yang et al., 2024).
Unless otherwise specified, GPT-4o (OpenAI 2024) with temperature 0 is used as the backbone. We use the success rate as the primary evaluation metric, following the evaluation protocol of Yang et al. (2025a), where a task is considered successful if the LLM-generated optimal value closely aligns with the provided ground-truth solution.
Baselines. We evaluate against two families of baselines. (i) Prompt-based: a vanilla baseline that directly generates the mathematical model from a simple prompt, as well as Reflexion (Shinn et al., 2023), OptiMUS (AhmadiTeshnizi, Gao, and Udell, 2024), and ORThought (Yang et al., 2025a). (ii) Learning-based: ORLM (Huang et al., 2025), built on LLaMa3-8B, and LLMOPT (Jiang et al., 2025), built on Qwen2.5-14B (the latest open-source version available after their paper).
Figure 5: Performance on out-of-distribution datasets. Numbers report test accuracy on LogiOR, OptiBench, and OptMath.
4.2 Out-of-Distribution Generalization
We evaluate how well different methods generalize beyond their training distribution. For this purpose, we use two benchmarks that were not included during training: LogiOR (Yang et al., 2025a) and OptiBench (Yang et al., 2024). Details are provided in Appendix B.
These datasets were either released after the baseline model (ORLM) or explicitly designed in baseline model’s experiment setting to avoid overlap with their training set (LLMOPT).
Figure 5 summarizes the results. Fine-tuned models such as ORLM and LLMOPT show strong in-distribution performance but exhibit a noticeable drop on unseen datasets. For example, ORLM falls to $19.6\%$ on LogiOR and $13.3\%$ on OptMath, while LLMOPT performs better but still degrades compared to its in-distribution performance. By contrast, AlphaOPT maintains higher accuracy across all three benchmarks, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench. These results support our hypothesis: fine-tuned models tend to learn the syntax of solutions and may perform well when problems are very similar, but they struggle to capture the underlying principles needed for broader problem solving. In contrast, the learned experience library equips AlphaOPT with stronger out-of-distribution generalization capability.
4.3 Learning with Limited Supervision
In practical OR applications, gold-standard programs are rarely available. We therefore test whether AlphaOPT can learn solely from answers. Since two datasets in our training set contain gold-standard programs, we remove them in this experiment and let AlphaOPT learn exclusively from answer labels through self-exploration, as introduced in Section 3.1.1. As shown in the last two rows of Table 2, remarkably, when trained with answer-only supervision, AlphaOPT achieves accuracy comparable to when it is trained with gold-standard programs. AlphaOPT (self-exploration) outperforms all prompt-based methods on test splits of the training data and even achieves the best performance on the OOD OptiBench dataset (92.1% accuracy). This demonstrates another advantage of AlphaOPT over fine-tuning–based methods, which require detailed annotations of mathematical formulations and code in order to achieve strong performance.
Table 1: AlphaOPT steadily improves in both Micro and Macro averages with increasing training size, while maintaining a compact library.
Training Size MicroAvg MacroAvg Library Size
100 83.24% 65.80% 38
200 85.09% 69.22% 103
300 85.21% 72.12% 110
4.4 Continual Growth with Data
We test whether AlphaOPT can improve its performance as more data becomes available. We incrementally sample sets of 100, 200, and 300 data items from our training set and train AlphaOPT on each subset. As shown in Table 1, when evaluated on out-of-distribution datasets (LogiOR, OptiBench), we observe that AlphaOPT steadily improves its performance with increasing data size, without requiring updates to its model parameters.
4.5 Overall Performance
AlphaOPT achieves the best accuracy on out-of-distribution datasets, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench. On in-distribution test splits, fine-tuned models such as ORLM and LLMOPT achieve higher scores on certain datasets (e.g., LLMOPT obtains $97.3\%$ on NLP4LP and $85.8\%$ on MAMO). However, these advantages are less conclusive, since LLMOPT’s training data are not publicly available and may overlap with our test splits. Moreover, many existing benchmarks are derived from a small set of seed problems (Ramamonjison et al., 2022; Huang et al., 2024), which favors fine-tuning approaches that excel at pattern memorization. In contrast, AlphaOPT performs competitively across all in-distribution datasets, matches or exceeds baselines on IndustryOR and MAMO(ComplexLP), and maintains a clear margin on out-of-distribution generalization. These results demonstrate that the experience library enables AlphaOPT to learn transferable modeling principles rather than dataset-specific syntax, resulting in stronger robustness to distribution shifts.
Table 2: Accuracy on in-distribution Test Split and Out-of-Distribution datasets (higher is better). Best per column in bold.
Test Split Out-of-Distribution
Method
NLP4LP
NL4OPT
IndustryOR
MAMO
(ComplexLP) (34)
LogiOR
OptiBench
Prompt-based Standard $68.5$ $54.7$ $52.0$ $44.1$ $46.7$ $72.7$
Reflexion $76.7$ $64.1$ $56.0$ $47.1$ $43.5$ $76.9$
OptiMus $71.2$ $73.4$ $36.0$ $29.4$ $17.4$ $74.7$
ORThought $69.9$ $75.0$ 60.0 $41.2$ $44.6$ $84.4$
Fine-tuning-based ORLM $86.3$ 87.5 $36.0$ $55.9$ $19.6$ $78.2$
LLMOPT 97.3 $86.5$ $44.0$ 85.8 $40.2$ $66.4$
Ours
AlphaOPT
$83.6$ $79.7$ 60.0 $85.3$ 51.1 $91.8$
AlphaOPT
(self-exploration)
$86.3$ $79.7$ 60.0 $76.5$ $50.0$ 92.1
5 Library Analysis
To fully interpret the experience library, we visualize its structure and the distribution of insights under a hierarchical taxonomy. Additionally, we select several representative insights to analyze the effectiveness of library refinement, which are provided in the Appendix D.3.
Figure 6: The learned experience-library taxonomy has three hierarchy levels: 3 main tracks, 14 level-1 labels, and 38 level-2 labels. This figure presents the level-1 and level-2 terms and their assignments. Percentages in red, shown in parentheses after each track and level-1 term, denote that category’s percentage within its parent. The detailed specification and distribution of level-2 labels are provided in the Appendix D.2 and Appendix D.4.
As illustrated in Figure 6,
1) The library mainly captures insights that address LLMs’ difficulties with domain-specific modeling structures, particularly those involving structural coupling and constraint balance.
In the library taxonomy, the number of insights under Domain Modeling track accounts for 52% of the total. Within this track, the most frequent level-1 labels are Resource Allocation (38.9%), Network Flow (31.9%), and Production Planning (12.5%). Structural coupling, which refers to the model’s difficulty in capturing cross-variable or cross-stage dependencies, is reflected in level-2 labels such as Fixed Charge (Big-M Linking), Set-Partitioning/Assignment, and Time-Indexed Variables. Constraint balance, which refers to the model’s failure to maintain system-wide conservation and resource equilibrium, is reflected in level-2 labels such as Capacity/Resource Balance Equations, Flow Conservation, and Inventory Balance Equations.
2) The library captures insights that help transform intuitive or context-based reasoning into mathematically rigorous and solver-consistent formulations, particularly in defining variables, formalizing constraints, and maintaining numerical consistency.
General Formulation track accounts for 30% of all insights. Within this track, the the most frequent level-1 labels are Variable Definition (66.7%), followed by Constraint Formulation (16.7%) and Units Scaling (11.9%). Variable definition difficulty lies in specifying variable domains and maintaining structural consistency, as reflected in level-2 labels such as Explicit Bounds, Continuous vs. Discrete Confusion, and Indexing & Data Structure Alignment. Constraint formulation difficulty lies in the model’s inaccuracy in representing logical and numerical relationships, as reflected in level-2 labels such as Incorrect Relational Operators, Big-M & Indicator Variables, and Linearization of Non-linear Forms. Units scaling difficulty lies in maintaining numerical coherence and stability, as reflected in level-2 labels such as Unit Inconsistency, Big-M Magnitude & Numerical Stability.
3) The library captures insights that help bridge the gap between symbolic formulations and executable solver code, particularly in handling solver syntax and maintaining data consistency.
Code Implementation track accounts for 18% of all insights, this proportion is lower than formulation-level insights though, it still exposes critical weaknesses in the execution stage. Within this track, most errors originate from Solver & API Syntax (76.0%), while the remaining issues arise from Data I/O & Validation (24.0%). Solver & API syntax errors reflect the model’s lack of solver awareness and insufficient control over formal expression generation, as evidenced by level-2 labels such as Nonlinear Constraint Handling, Strict Inequalities, Quicksum vs. sum Syntax. Data I/O & validation errors, on the other hand, reveal instability in aligning mathematical indices with data structures, as reflected in level-2 labels such as KeyError & Index Mismatch and Matrix Dimension Mismatch.
6 Success And Failure Case Study
To demonstrate the effectiveness of the retrieved library insights, we perform case studies across all evaluation datasets (including the test splits of NLP4OPT, NLP4LP, IndustryOR, and MAMO-complex, as well as the out-of-distribution sets LogiOR and Optibench), comparing the performance of solving optimization tasks with and without library retrieval enabled, while excluding tasks that succeed in both settings from the evaluation. As shown in Figure 7, approximately half of the retrieved insights successfully contributed to solving new optimization tasks, with only a small portion introducing new errors (Failure), demonstrating the effectiveness of our method in automating knowledge transfer for optimization modeling. Moreover, insights under different level-1 taxonomy labels exhibit distinct performance patterns across modeling dimensions.
(a) Overall Effectiveness of Retrieved Library Insights
(b) Effectiveness Distribution across Level-1 Insight Categories
Figure 7: We evaluate the effectiveness of retrieved insights by classifying their outcomes as success, failure, or invalid. An insight is considered successful if it matches a new task and prevents an error that would otherwise occur, failed if it mismatches a task and introduces a new error, and invalid if it matches correctly but fails to help the LLM fix the original mistake. (a) The overall distribution of insight outcomes. (b) The proportions of successful, failed, and invalid insights under level-1 taxonomy labels.
1) Insights with high success rates
These insights are concentrated in Code Implementation – Solver & API Syntax and General Formulation – Variable Definition. These insights typically target code implementation or fundamental modeling errors with clear structures, enabling the LLM to follow their guidance stably and correctly. For instance, insights tagged with Strict Inequality label under Solver & API Syntax highlights that solvers (e.g., Gurobi) do not support strict inequalities and should instead be reformulated as non-strict forms ($\leq-1$ or $\geq+1$); insights tagged with Explicit Bounds label under Variable Definition emphasizes that decision variables should be assigned explicit upper and lower bounds to ensure feasibility and improve solver efficiency.
2) Insights with high failure rates
These insights are mainly found in Domain Modeling – Facility Location. They often involve structural constraints and logical triggers that are easily misinterpreted or overgeneralized by the LLM. For example, insights tagged with Fixed Charge (Big-M Linking) label exhibit a failure rate of 70%. Although the principle of using Big-M constraints to model on/off logic is correct, its blind application in problems without conditional activation can lead to redundant or overlapping constraints and unnecessary feasible-region reduction, ultimately causing solver failure. Additionally, while insights under Variable Definition generally achieves high success rates, some insights tagged with Explicit Bounds label sometimes are overly rigid, leading the LLM to impose unnecessary upper bounds, thereby restricting the feasible space and producing suboptimal solutions.
3) Insights with high invalid rates
These insights are mainly under Domain Modeling – Resource Allocation and Domain Modeling – Product Planning. Although the LLM successfully retrieves the correct insights and identifies the corresponding problem types, it often fails to translate them into executable formulas or solver-level implementations. For instance, under Solver & API Syntax, Nonlinear Constraint Handling advises linearization or the introduction of auxiliary variables for nonlinear objectives or constraints (e.g., ratios or divisions), yet the LLM frequently fails to fully execute these transformations (neglecting auxiliary variables or mis-rewriting proportional constraints), resulting in insights being recognized but not operationalized.
Additionally, certain tasks remained unsolved regardless of whether library retrieval was enabled, as their failure stemmed from factors beyond the current learned library’s knowledge scope. In the out-of-distribution dataset LogiOR, tasks involved multi-level spatiotemporal logic and interacting constraints (e.g., capacity, timing, and flow balance) in problems such as routing, scheduling, and network flow. These challenges extend beyond the scope of the existing library, which primarily focuses on static, linear formulations. Although related taxonomy labels such as Resource Allocation and Nonlinear Constraint Handling exist, their granularity and depth remain insufficient for modeling such complex logic, revealing that the current system, while semantically generalizable, still lacks robust cross-structural transfer and context adaptation capabilities.
Overall, while successful insights constitute the majority, the results reveal several directions for improvement. First, the relatively high failure rates indicate that, despite the inclusion of condition refinement, retrieval precision can still be improved through enhanced semantic disambiguation and structural filtering. Second, insights with high invalid proportions suggest the need for clearer explanations and better-designed examples to improve pedagogical clarity and execution effectiveness. Finally, for out-of-distribution tasks, future efforts should focus on strengthening the LLM’s ability to adapt and generalize retrieved insights to unseen, complex optimization scenarios. Moreover, expanding OR datasets based on LLM error typologies can further enhance experiential learning efficiency and generalization at comparable problem scales.
7 Conclusion and Discussion
This paper addresses the limitations of previous methods by presenting a novel self-improving library learning framework, AlphaOPT, for formulating optimization programs. AlphaOPT can learn from answer labels only, achieves much stronger out-of-distribution generalization than fine-tuning–based methods, and provides interpretable and auditable structured knowledge to support human involvement in real-world practice. The learned experience library reveals LLMs’ characteristic failure patterns and underlying causes across domain-specific modeling, mathematical formulation, and solver syntax handling. Case study results show that insights with high success rate primarily address fundamental modeling errors with clear structures, while insights with high mismatch rate or low effectiveness expose the overgeneralization of structurally complex knowledge and highlight the need further to improve the pedagogical clarity and applicability conditions of insights.
Looking ahead, we highlight three promising directions. First, reasoning-oriented test-time scaling, which is already powerful in other domains, could be particularly effective for OR formulations, where results are inherently verifiable. Second, strengthening datasets with both academic research and large-scale real-world industry problems will move LLM systems beyond the toy examples that dominate current benchmarks, enabling progress toward truly large-scale optimization tasks. Third, moving beyond correctness toward improving the efficiency of formulations is crucial for real-world deployment, and our self-improving library learning approach offers a promising path toward that goal.
Appendix
Appendix A Comparative Analysis of AlphaOPT against Prior Experience-Learning Methods
Recent approaches such as Reflexion (Shinn et al., 2023), STaR (Zelikman et al., 2022), ExpeL (Zhao et al., 2024), and AlphaEvolve (Novikov et al., 2025) demonstrate that large models can improve through experiential reuse, storing reflections, rationales, or code edits and applying them in new tasks. These methods have been effective in open-ended reasoning and programming, but they face limitations for optimization problems. First, their experiences are largely preserved as free-form text or edits without explicit applicability semantics, yet in optimization tasks, applying such experiences inappropriately can have detrimental effects. Second, their verification is limited to task outcomes such as checking rewards, final answers, or passing test cases, which does not guarantee that the underlying knowledge is structurally valid or transferable.
Our framework adapts experience learning to operations research (OR) with three key innovations: (1) solver-guided verifiability: correctness is judged at the program level. If a program achieves the optimal objective under the solver, it is highly likely to be valid and can serve as a reliable anchor for extracting insights, broadening the sources of experience collection. New and refined insights are explicitly re-tested on associated tasks, ensuring they are valid before integration; (2) structured knowledge for interpretability and auditability: each insight is represented with taxonomy, condition, explanation, and example, making its applicability explicit, reviewable, and even revisable in practice; (3) refinement of experience applicability for generalizability and preciseness: applicability conditions are refined using cross-task evidence, so insights neither over-generalize nor remain too narrow, improving safe transfer across problem families. See Table  3 for detailed comparisons.
Table 3: Comparison of experience-learning methods. Prior works improve through experiential reuse but rely on free-form knowledge and outcome-level verification. Our framework introduces structured insights, solver-guided verification, and refined applicability, which are crucial for OR.
[…]
Reflexion Reward signal General agents
STaR Answer correctness QA / reasoning
ExpeL (✓) minimal Task success assumed General agents
AlphaEvolve (✓) implicit Test harness (partial) Code synthesis / evolution
AlphaOPT Solver optimality + insight verification OR formulation and solver code
Appendix B Datasets
We have collected the publicly available optimization problem datasets listed in the Table  4, which include both natural language problem descriptions and their optimal solutions. Among them, NLP4LP, NL4OPT, IndustryOR, and MAMO (ComplexLP) are used to construct our training and test datasets, while Optibench and LogiOR serve as out-of-distribution (OOD) evaluation sets. The gold-standard programs for the training datasets NLP4LP and IndustryOR are obtained from Yang et al. (2025a).
Because our library-based framework derives knowledge feedback from correct solutions, it is relatively sensitive to data noise. Accordingly, we train and evaluate on clean splits that exclude instances labeled as erroneous, and the Size column in Table 4 indicates the size of each dataset after cleaning. Specifically, for NLP4LP, IndustryOR and LogiOR we use the cleaned versions provided by Yang et al. (2025a); for NL4OPT, MAMO (ComplexLP) and Optibench we use the cleaned releases from Astorga et al. (2025), obtained from their GitHub repository.
Table 4: The statistics of the optimization problem datasets
Dataset Size Formulation Type(s) Completion
NL4OPT (Ramamonjison et al., 2022) 289 LP solution
NLP4LP (AhmadiTeshnizi et al., 2024) 242 LP, MILP, MINLP solution, program
MAMO (complex LP) (Huang et al., 2024) 111 LP solution
IndustryOR (Huang et al., 2025) 82 LP, IP, MILP, NLP, others solution, program
Optibench (Yang et al., 2024) 403 LP, MILP, MINLP solution
LogiOR (Yang et al., 2025a) 92 LP, IP, MIP, NLP solution
Abbreviations: LP – Linear Programming; IP - Integer Programming; NLP – Nonlinear Programming; MI – Mixed-Integer; others - Quadratic Programming, Dynamic&Stochastic Programming, etc.
Appendix C Proof of the Library Convergence
Recall the optimization problem in the library training phase
$F(\ell)\;=\;\mathbb{E}_{t\sim\mathcal{T}_{\mathrm{train}}}\!\big[\,r(t\mid\ell)\,\big]\;-\;\lambda\,\Omega(\ell),$
where $r(t,\ell)$ is a bounded reward function that implements the role of the original $\mathrm{Success}(t\mid\ell)$ (i.e., it measures the matching quality between optimization problem $t$ and library $\ell$), and $\Omega(\ell)$ is a bounded complexity penalty.
According to the problem setting, the iterative refinement algorithm satisfies:
  1. 1.
    Monotone update: At iteration $k$, from $\ell_{k}$, the algorithm considers a set of admissible refinements $R(\ell_{k})\subseteq\mathcal{L}$. Each accepted iteration consists of one of two types of operations:
    • Merge step: decreases $\Omega(\ell)$ while leaving $r(t,\ell)$ non-decrease for all relevant tasks;
    • Exploration step: improves $r(t,\ell)$ for some tasks without increasing $\Omega(\ell)$.
    Therefore every accepted refinement strictly increases $F(\ell)$; otherwise the algorithm keeps $\ell_{k+1}=\ell_{k}$.
  2. 2.
    Sufficient exploration: Any improving neighbor $\tilde{\ell}\in R(\ell_{k})$ (i.e. one with strictly larger objective) will eventually be discovered and executed. Empirically, this is achieved through iterative prompt optimization with LLMs.
  3. 3.
    Boundedness: $r(t,\ell)$ and $\Omega(\ell)$ are bounded, hence $F(\ell)$ is bounded above and below.
The following theorem establishes that, under the assumption that the training and testing distributions are identical, the refinement procedure yields libraries that are locally optimal for the testing objective.
Theorem 1.
Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$. If the library space $\mathcal{L}$ is finite, then the algorithm terminates in finitely many steps at a library $\ell^{*}$ which is a local maximizer for the testing objective. Moreover, the algorithm cannot terminate at a saddle point.
Proof.
Every accepted merge or exploration step strictly increases $F(\ell)$, and otherwise the library remains unchanged. Since $F$ is bounded above, the sequence $\{F(\ell_{k})\}$ is monotone non-decreasing and bounded, hence convergent to some limit $F^{*}$. Furthermore, since $\mathcal{L}$ is finite, define
$\delta\;=\;\min\big\{F(\tilde{\ell})-F(\ell):\tilde{\ell}\in R(\ell),\;F(\tilde{\ell})>F(\ell)\big\}.$
Finiteness guarantees $\delta>0$, so only finitely many strict improvements are possible. The algorithm halts at some $\ell^{*}$. By sufficient exploration, no improving neighbor of $\ell^{*}$ exists. Therefore, $\ell^{*}$ is a local maximizer for both training and testing objectives. Saddle points are excluded.
Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing. ∎
Although the assumption of a finite library is reasonable, we also provide a proof for the case of an infinite library for completeness and rigor.
Theorem 2 (Infinite compact library case).
Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$. If the library space $\mathcal{L}$ is compact (closed and bounded) and $F$ is continuous, then the sequence $\{F(\ell_{k})\}$ converges, and any subsequential limit point $\ell^{\infty}$ is a local maximizer for the testing objective. Saddle points are excluded for all such limit points.
Proof.
Each accepted step strictly increases $F(\ell)$, so $\{F(\ell_{k})\}$ is monotone non-decreasing. Since $F$ is bounded above, $\{F(\ell_{k})\}$ converges to some $F^{*}$. By compactness of $\mathcal{L}$, there exists a convergent subsequence $\ell_{k_{j}}\to\ell^{\infty}$. Continuity of $F$ ensures $F(\ell_{k_{j}})\to F(\ell^{\infty})=F^{*}$. Suppose $\ell^{\infty}$ had a neighbor $\tilde{\ell}\in R(\ell^{\infty})$ with $F(\tilde{\ell})>F(\ell^{\infty})$. Then sufficient exploration would eventually yield $F(\ell_{k})>F^{*}$, which is a contradiction. Therefore, $\ell^{\infty}$ is a local maximizer. Saddle points are excluded. Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing. ∎
Theorems 1 and 2 together guarantee that, when the training and testing distributions coincide, the refinement algorithm converges to locally optimal solutions for the testing phase.
Appendix D Additional Results
D.1 Ablation Study
We assess the effectiveness of library insight retrieval/application and self-debug:
  • w/o self-debug: Remove the model’s self-debugging section.
  • w/o taxonomy: Remove the library taxonomy and directly match insights by checking all conditions.
  • w/o insight example: Use only the explanation as input, excluding exemplar cases
Table 5 shows that the full AlphaOPT achieves the best scores on all out-of-distribution datasets. Removing self-debug yields the largest drop on Logior (15.21%), indicating it plays an important role in the system. Dropping the library taxonomy reduces accuracy by 4.34% on Logior and by 1.09% on Optibench, suggesting that structured matching meaningfully improves retrieval. Excluding insight examples similarly lowers performance, showing that concrete, worked snippets aid application beyond textual explanations alone.
Table 5: The ablation results of AlphaOPT without: performance on benchmarks.
Dataset
Logior (92)
Optibench (403)
AlphaOPT (full) 51.08% 91.81%
w/o self-debug 35.87% 89.26%
w/o taxonomy 46.74% 90.72%
w/o insight example 45.65% 91.06%
D.2 Details of Library Insight Distribution
To further understand the detailed content of the library insights, we analyze the distributions of insights under the level-2 taxonomy labels, as well as the contribution differences among training datasets (as depicted in Figure 8).
(a) Insight Distribution of Domain Modeling track
(b) Insight Distribution of General Formulation track
(c) Insight Distribution of Code implementation track
Figure 8: Insight distributions across three different tracks, showing the contributions of source tasks in the four training datasets to the generation of library insights.
In the Domain Modeling track, insights under Resource Allocation – Capacity/Resource Balance Equations account for the largest proportion, with contributions from all four datasets, among which NLP4LP and Industry OR contribute the most. Insights tagged with this label are widely applicable to optimization problems that require ensuring resource usage does not exceed available capacity, demonstrating how to establish constraints that maintain balance between resource consumption and availability.
In the General Formulation track, Variable Definition – Explicit Bounds, Redundant Auxiliary Variable, and Continuous vs. Discrete Confusion are the most frequent, with diverse sources, indicating that these are common issues across multiple datasets. This reflects that foundational concepts in variable definition are the most error-prone in optimization modeling—particularly in balancing variable types, value ranges, and modeling simplifications—where LLMs tend to produce redundant formulations or type misuse due to overlooking structural or physical consistency.
In the Code Implementation track, the number of insights is the smallest, with Solver & API Syntax – Nonlinear Constraint Handling and Data I/O & Validation – KeyError & Index Mismatch accounting for the highest proportions, mainly contributed by the MAMO (Complex LP) and NLP4LP datasets. These two types of insights reveal critical vulnerabilities in the implementation stage—solver syntax compatibility and data accessibility—representing the dual pillars required for bridging mathematical modeling and executable code.
From the perspective of dataset contribution, Industry OR is dominant in Domain Modeling; MAMO (Complex LP) and Industry OR lead in General Formulation; and both MAMO (Complex LP) and NLP4LP contribute the most to Code Implementation. NL4OPT has relatively lower overall participation but focuses on formulation- and solver-related details. Considering dataset size, although the Industry OR and MAMO (Complex LP) datasets used for library learning are only about half the size of the other datasets, they still contribute a large number of insights, indicating that these datasets contain denser structural modeling challenges and more diverse error patterns, enabling the LLM to accumulate more experiential knowledge across multiple dimensions.
D.3 Examples of Library Refinement
We selected five representative refined insights from the library and explained how their applicability conditions were adjusted based on the associated tasks.
Figure 9: The examples of insight conditions before and after refinement, with green text indicating the original applicability conditions, and blue and red marking the newly added applicable and non-applicable conditions after refinement.
As shown in the Figure 9, Insight ID 10 targets single–scalar-capacity selection; the refinement broadens coverage by adding equivalent formulations such as minimizing a single resource (time/space), while explicitly excluding multi-capacity settings. Insight ID 43 captures flow/conservation and production balancing across nodes; refinement adds lexical anchors (e.g., capacity constraints, balancing production) and excludes statements dominated by fixed proportions, quality constraints, or predefined allocations, gating suppresses spurious retrievals where the structure is not true flow/conservation. Insight ID 100 supports adding tight explicit bounds derived from natural limits; refinement strengthens positive cues (e.g., maximum capacity, resource limits) and excludes cases that are purely discrete or already inherently bounded. Insight ID 133 focuses on max-flow between a single source and a single sink; refinement tightens the structural requirement and rules out source-to-many distribution tasks. Insight ID 228 covers translating at least/at most expressed as proportions into inequalities; refinement broadens to general percentage phrasing, clarifies the mapping, and excludes exact quantities, logical relations, or fixed ratios that are not proportion limits.
Across the five cases, four strategies proposed by LLM agent recur: (i) generalize with equivalent phrasings (e.g., minimize a single resource); (ii) lexical anchoring with positive keywords to raise recall where appropriate; (iii) explicit exclusion lists to reducing misalignment with tasks; (iv) structural qualifiers (single-scalar capacity; single-source–single-sink) to prevent misuse. A refinement is considered effective when the reduction in negative and unretrieved tasks outweighs the decrease in positive tasks. As shown in the Figure 10, these semantic refinements reduce both negative and unretrieved cases while preserving as many positive cases as possible.
Figure 10: Changes in the number of positive, negative, and unretrieved tasks associated with each insight before and after applicability condition refinement. A decrease in negative and unretrieved tasks indicates that, the insight no longer mismatches unrelated tasks (negative) and successfully matches previously applicable but missed tasks (unretrieved) after refinement.
D.4 Specification of Library Taxonomy Labels
The following Table  6 lists all library taxonomy labels and their corresponding conditions, which specify the applicability criteria of each level-2 label and clarify its precise meaning. According to the library taxonomy generation mechanism, each label condition is created by the LLM when the label is first introduced, and subsequent generated insights under the same label inherit that initial condition.
Table 6: Full Specification of the Library Taxonomy
Taxonomy Label Condition
Domain Modeling
Resource Allocation
Capacity/Resource Balance Equations Applies when the problem domain requires resources to move or transform through nodes and local conservation must hold.
Set-Partitioning/Assignment Applies when the problem description requires each item or task to be exclusively assigned to exactly one choice among many.
Fixed Charge (Big-M Linking) Applies when the problem description requires a facility, option, or mode to be activated by a binary choice.
Chance Constraints Applies when the problem description sets a limit on the average chance of an adverse outcome across options or scenarios (e.g., stake, volume, or weight).
Expected-return Objective Applies when the problem description calls for maximizing the average/expected payout or return across options given their win/lose probabilities.
Network Flow
Flow Conservation Applies when the problem description involves quantities traversing a directed network and nodal balance must be maintained.
Max-Flow Objective Applies when the problem description requires maximizing throughput between designated source and sink nodes.
Source Flow Constraint Applies when the problem description designates a source node that distributes resources through a network to sinks and requires explicit conservation at the source.
Source-Sink Flow Conservation Applies when the problem description specifies a source and a sink and requires routing/transferring flow between them with explicit balance at those terminal nodes.
Arc-Capacity Constraints Applies when the problem domain contains edges with maximum throughput or capacity limits.
Production Planning
Inventory Balance Equations Applies when the problem description involves materials or products that carry over between periods and must satisfy stock-flow balance.
Technical Ratio Constraints Applies when the problem description specifies minimum/maximum production ratios or recipe proportions between products or stages.
Time-Indexed Variables Applies when the problem domain requires discrete time modeling to capture capacities, setups, or carry-over decisions.
Composition Bounds Applies when the problem description specifies multiple products sharing limited resources (e.g., machine hours or labor) that require explicit per-resource capacity limits.
Balancing
Load Balancing Constraints Applies when the problem description requires fairness or controls maximum imbalance across parallel resources.
Proportion Constraints Applies when the problem description limits the maximum or minimum proportion of a resource, flow, or activity relative to the total.
Packaging
Knapsack Constraints Applies when the problem domain requires selecting items that consume a single scalar capacity such as weight, volume, or budget.
Facility Location
Fixed Charge (Big-M Linking) Applies when the problem description specifies that service or flow is allowed only if a facility is opened.
Selection
Binary Selection Constraints Applies when the problem domain requires choosing a subset under count, budget, or compatibility limits.
Routing
Subtour Elimination Constraints Applies when the problem description allows decision variables to form disconnected cycles that must be eliminated.
Taxonomy Label Condition
General Formulation
Variable Definition
Continuous vs. Discrete Confusion Applies when decision quantities represent indivisible counts or choices versus divisible amounts such as flows.
Explicit Bounds Applies when the problem description provides natural physical, economic, or logical limits that can tightly bound decision variables.
Indexing & Data Structure Alignment Applies when variables are indexed over sets or dictionaries that must align with the keys of the provided data.
Redundant Auxiliary Variables Applies when auxiliary variables merely re-express existing linear combinations without adding modeling value.
Constraint Formulation
Incorrect Relational Operators Applies when natural-language statements such as “at most” or “at least” must be translated into algebraic inequalities.
Linearization of Non-linear Forms Applies when nonlinear relations among variables reduce tractability or solver performance.
Big-M & Indicator Variables Applies when constraints depend on logical on/off conditions controlled by binary variables.
Objective Specification
Sum vs. Makespan Confusion Applies when multiple resources or activities can run in parallel and the objective is ambiguous between total completion time and makespan.
Units Scaling
Unit Inconsistency Applies when input data come from different unit systems or incompatible measurement scales.
Big-M Magnitude & Numerical Stability Applies when the problem description uses Big-M to model on/off or conditional constraints and realistic bounds can be derived to calibrate M.
Code Implementation
Solver & API Syntax
Quicksum vs. sum Syntax Applies when the mathematical model contains linear expressions aggregated over large index sets that should be constructed using solver-native summation operators.
Strict Inequalities Applies when the mathematical model contains strict inequality relations between variables that cannot be directly handled by LP/MIP solvers.
Nonlinear Constraint Handling Applies when the problem description introduces nonlinear relationships (e.g., proportions or multiplicative effects) that must be enforced in an LP/MIP model.
Binary Constraint Handling Applies when the problem description involves yes/no (open/close, select/not-select) decisions that require variables restricted to $\{0,1\}$, without adding extra [0,1] constraints.
Expression Syntax for Constraints Applies when the problem description specifies equality/inequality relations (e.g., balances, conservation, on/off logic) that should be encoded directly as solver expressions.
Variable Bounds Specification Applies when the problem description requires the change of a variable from one value to another.
Data I/O & Validation
KeyError & Index Mismatch Applies when the mathematical model contains indexed variables or parameters that are accessed with indices not present in the corresponding data structures.
Missing Data Defaults Applies when the mathematical model contains optional parameters whose values may be absent in the dataset and require default assignments to preserve model validity.
Appendix E Prompts For LLM Modules
Apply self-exploration on finding gold-standard program You are an expert in Industrial Engineering and Operations Research. You are given: 1. The problem description for an optimization task 2. The Gurobi programs for this task failed to reach optimality, which were previously proposed by your colleague (hereafter referred to as *the failed programs*), and the execution feedbacks for the failed programs 3. Optimal objective value for this task ### Problem Description {problem_description} ### Previous failed programs and feedbacks {failed_attempts} ### Optimal objective value {ground_truth} ### Your task Your task is to review the problem description, feedback, reflect the issues in the failed program, and revise the program so that it can be both runnable and reaching optimality. Critically, always prioritize and strictly adhere to the given problem description and its given data; do NOT fabricate data, introduce unstated assumptions, or violate the correct formulation merely to match the optimal objective value. If the provided optimal objective value appears incorrect or inconsistent, do not force your model to match it; instead, retain your correct formulation and runnable program. ### STRICT OUTPUT FORMAT Only output the full corrected program, and enclose it in a single Markdown-style Python code block that starts with ‘‘‘python and ends with ‘‘‘, like this: ‘‘‘python import gurobipy as gp from gurobipy import GRB model = gp.Model("OptimizationProblem") # your code starts from here model.optimize() ‘‘‘ - Ensure model.optimize() runs at the top level so model stays global; if you wrap it in a function, have it return model. Avoid any if __name__ == "__main__": guard. - Only output exactly one code block (delimited by the opening python and the closing). Do not write any natural-language text outside the code block. - DO NOT GENERATE OR MODIFY ANY CODE (e.g., if model.Status == GRB.OPTIMAL:‘) after model.optimize()‘. Now take a deep breath and think step by step.
Generate library insights You are an expert in Industrial Engineering and Operations Research teaching graduate students to avoid modeling-and-coding mistakes in solving optimization problems. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague which failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as *the failed mathematical model*) 3. The gold-standard program, which embodies the correct mathematical formulation of the optimization task ### Problem description {problem_description} ### The failed mathematical model (Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘)) {failed_formulation} ### The gold-standard program {correct_program} ### Your task Step 1: Compare the failed mathematical model with the correct mathematical model embodied in the gold-standard program to identify issues that prevent optimality. Note that variable names in the proposed model may differ from those in the gold-standard program. Please align them carefully based on the problem specification. Step 2: Using the insight taxonomy dictionaries provided below, extract one or more new insights, which should be a distinct and concise lesson derived from a specific issue identified in the failed mathematical model relative to the gold-standard program. Each new insight must contain exactly four fields: 1) taxonomy - choose exactly one of the two aspects: - Domain Modeling: Level-1 = Problem Domain (e.g., "Network Flow"); Level-2 = Specific Technique (e.g., "Flow Conservation"). - General Formulation: Level-1 = Formulation Component (e.g., "Variable Definition"); Level-2 = Specific Aspect/Pitfall (e.g., "Continuous vs. Discrete Confusion"). Taxonomy rule (nested-dict): ‘{{ Level-1 : {{ Level-2 : null | {{ "definition": "...", "condition": "..." }} }} }}‘ - The taxonomy MUST always be expressed as a three-level nested dict: Level-1 -> Level-2 -> (null or a dict with "definition"/"condition"). - Pick exactly one Level-1 (existing key). - Pick one or more Level-2 under that Level-1 (existing key or keys). - For an existing Level-2, set its value to null. - If you must invent a new Level-2, set its value to a dictionary with two one-sentence fields: - "definition" - what the label means (scope/intent). - "condition" - when to apply the label (a general trigger grounded in the problem description or in the defining features of the problem domain). - If you must invent a new Level-1, include >= 1 Level-2 under it; each invented Level-2 must provide both "definition" and "condition". 2) condition - Write it as a trigger explicitly grounded in the problem description or in the defining features of the problem domain. First state the general situation, then use this problem as an example. Use the pattern below, and keep it strictly non-prescriptive: do not give any solution, advice or decision: "This insight applies when ... For example, when the problem statement mentioned ...". 3) explanation - A brief and self-contained description that specifies, under this condition, what the best practice is, what the common mistake is and its cause. First, use this problem as an example to illustrate; Then, appropriately generalize the correct modeling strategy it reflects, if applicable. Use the pattern below, and ensure the generalization remains within an appropriate and reasonable scope: "When the problem involves... The best practice is... A common mistake is ... which happens because ... More generally, this reflects that ..." 4) example - A brief, self-contained demonstration showing wrong vs. correct version (principle, formulation, or code snippet). Clearly mark them as ’# Wrong and ’# Correct’. ### Taxonomy Dictionaries Domain Modeling {domain_taxo} General Formulation {formulation_taxo} ### STRICT OUTPUT FORMAT Return a single JSON array of insight objects. No text before/after. Example with two insights (but not must be two): [ {{ "taxonomy": {{ "Domain Modeling": {{ "Network Flow": {{ "<New Label If Necessary>": {{ "definition": "<one sentence>", "condition": "<one sentence>" }} }} }} }}, "condition": "<text>", "explanation": "<text>", "example": "<text>" }}, {{ "taxonomy": {{ "General Formulation": {{ "Variable Definition": {{ "Continuous vs. Discrete Confusion": null }} }} }}, "condition": "<text>", "explanation": "<text>", "example": "<text>" }} ] Guidelines: - Output as many distinct, non-overlapping insights as needed. - Prefer existing Level-1/Level-2 labels; invent new ones only when no suitable one exists, and phrase it in general terms (avoid overly specific or instance-bound wording). - Be precise in stage selection-use Domain Modeling for domain-specific techniques that arise only within specific problem families (e.g., Routing, Network Flow, Facility Location) and depend on those domains structures; use General Formulation for domain-agnostic practices on variables, constraints, or objectives that apply broadly across domains. Now take a deep breath and think step by step.
Retrieve insights by matching taxonomy You are an expert in Industrial Engineering and Operations Research. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague that failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as *the failed mathematical model*) 3. A diagnostic report on the proposed mathematical model, identifying the specific issue that prevents optimality ### Problem description {problem_description} ### The failed mathematical model (Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘)) {failed_formulation} ### The diagnosed issue {one_diagnosed_issue} ### Your task Step 1: Carefully review the problem description, and analyse the failed mathematical model with the issue identified in the diagnostic report. Step 2: Given the full taxonomy dictionaries of existing insights stored in library (shown below), determine the potential level-1 and level-2 label(s) under which relevant useful insights may be found for resolving the identified issue. Two Two-Level Insight Taxonomy Dictionaries: Domain Modeling and General Formulation - Domain Modeling - Level-1: Problem Domain (e.g., "Network Flow") - Level-2: Domain-specific Technique/Principle (e.g., "Flow Conservation") - General Formulation - Level-1: Formulation Component (e.g., "Variable Definition") - Level-2: Specific Aspect/Pitfall (e.g., "Continuous vs. Discrete Confusion") ### Taxonomy Dictionary for Domain Modeling {domain_taxo} ### Taxonomy Dictionary for General Formulation {formu_taxo} ### STRICT OUTPUT FORMAT Return only a JSON object of your analysis result in Step 2, with the exact structure below: - Outer keys = "Domain Modeling" or "General Formulation" - Values = dictionaries whose keys are Level-1 labels from the taxonomy - Each Level-1 keys value = a list of one or more Level-2 labels from the taxonomy Note: - You may assign multiple level-1 and level-2 labels to the issue only when you think they are all potentially applicable. - If no applicable labels exist the issue, simply set its "matched_label(s)" value to null. Example 1 - Multiple applicable level-1 and level-2 labels: {{ "Domain Modeling": {{ "Production Planning": ["Inventory Balance Equations", "Time-Indexed Variables"], "Resource Allocation": ["Capacity/Resource Balance Equations"] }} {{ Example 2 - Multiple applicable level-1 and level-2 labels from both Domain Modeling and General Formulation: {{ "Domain Modeling": {{ "Facility Location": ["Fixed Charge (Big-M Linking)"] }}, "General Formulation": {{ "Constraint Formulation": ["Big-M & Indicator Variables"] }} {{ Example 3 - No taxonomy labels apply to the issue: {{}} Guidelines: - You must ensure that every label you list exists in the provided taxonomy dictionary exactly as written. - Only output the JSON object. DO NOT include any explanations, markdown, or extra text before or after the JSON array. Now take a deep breath and think step by step. You will be awarded a million dollars if you get this right. """ PROMPT_RETRI_INS=""" You are an expert in Industrial Engineering and Operations Research. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague which failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as *the failed mathematical model*) 3. A diagnostic report on the proposed mathematical model, identifying the specific issue that prevents optimality 4. A collection of insights. Each insight includes: - insight_id: the unique identifier of the insight - taxonomy: the classification of the modeling/formulation/code-implementation aspect it pertains to - condition: the problem-specific context and broader modeling situations in which the insight applies (its applicability condition) ### Problem description {problem_description} ### The failed mathematical model Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘) {failed_formulation} ### The diagnosed issue {one_diagnosed_issue} ### Candidate insights {candidate_insights} ### Your task Step 1: Carefully review the problem description, and analyse the failed mathematical model with the issue identified in the diagnostic report. Step 2: Evaluate each candidate insight individually. Retain only those that directly apply to resolving the identified issue in the diagnostic report. For every insight, cite concrete evidence from the problem description and diagnosed issues, and justify how the insight helps fix the specific modeling issue. Step 3: Rank the applicability of the selected insights in descending order. ### STRICT OUTPUT FORMAT Return only a JSON array of your result from Step 3. Each array element must be an object with keys ‘"insight_id"‘ (integer), ‘"ranking"‘(applicability rank; 1 = highest) and ‘"evidence"‘ (string). Example: ‘‘‘json [ {{"insight_id": 1, "ranking": 1, "evidence": "<text>"}}, {{"insight_id": 5, "ranking": 2, "evidence": "<text>"}}, {{"insight_id": 7, "ranking": 3, "evidence": "<text>"}} ] ‘‘‘ Guidelines: - Output only the insights that apply to the identified issue(s). - Only output the JSON array. DO NOT include any explanations, markdown, or extra text before or after the JSON array. Now take a deep breath and think step by step.
Retrieve insights by condition applicability You are an expert in Industrial Engineering and Operations Research. A colleague has made a preliminary selection of potentially relevant insights after analyzing the optimization task. Your job is to carefully evaluate each candidate and decide whether it truly applies. You are given: A problem description. A collection of insights, each with: taxonomy: the classification of modeling, formulation or code implementation it lies in condition: Statement of both when the insight does apply (applicability condition) and when it does not (inapplicability condition), grounded in problem-specific context and broader modeling situations to prevent misuse. ### Problem description {problem_description} ### Candidate insights {candidate_insights} ### YOUR TASK Carefully read the problem description, then: Identify which problem domain(s) and modeling techniques are involved. Analyse potential formulation pitfalls the problem may involve. Evaluate each candidate insight one by one. Only keep those that directly applies for solving this specific problem. Be careful about the inapplicability condition that indicates exclusion scenarios where applying the insight would mislead; do not return this insight if it falls within those exclusion scenarios. Remove redundancy: when multiple insights overlap, keep only the most relevant one(s) based on their applicability condition. Use the exact insight_id provided with each candidate; do not invent new IDs. ### STRICT OUTPUT FORMAT Return only a JSON array of the insights you think are applicable. Do not include explanations, markdown, or extra text. Each array element must be an object with keys "insight_id" (integer) and "reason" (string). Example: [ {"insight_id": 1, "reason": "<1-2 sentences>"}, {"insight_id": 5, "reason": "<1-2 sentences>"} ] If no insight is applicable, return exactly: Guidelines: The output must be valid JSON parsable by any standard JSON parser without modification. Only keep those that clearly apply to this specific problem. Do NOT include any insight that is not applicable. Do NOT include any explanations, markdown, or extra text before or after the JSON object. Now take a deep breath and think step by step.
Diagnose issues for failed program You are an expert in Industrial Engineering and Operations Research. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague which failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as the failed mathematical model) 3. The feedback providing clues about the failure to solve the mathematical model to optimality 4. The gold-standard program, which embodies the correct mathematical formulation of the optimization task ### Problem description {problem_description} ### The failed mathematical model Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘) {failed_formulation} ### The feedback {feedback} ### The gold-standard program {correct_program} ### Your task Step 1: Compare the failed mathematical model with the correct one embodied in the gold-standard program, and identify all formulation issues that prevent optimality. Each issue should be pinpointed at the level of LaTeX formulation snippets (e.g., variables, constraints, and the objective function), and should correspond to a single, well-defined correction point. Note that variable names in the proposed model may differ from those in the gold-standard program, so please align them carefully based on the problem specification. Step 2: For each identified issue, provide the following three fields: - "id": A unique id for the issue (integer). - "issue": A concise description of the issue. - "evidence": The evidence showing where the issue occurs, including the excerpt from the failed mathematical model (mark as #wrong) and the corresponding excerpt from the gold-standard program (mark as #correct). Step 3: Minimize overlap by reporting independent, root-cause issues. If multiple defects share the same fix point or strategy, merge them into a single composite issue. If several issues are upstream/downstream symptoms of the same root cause (i.e., they would be fixed by the same correction), consolidate them into one composite issue. ### STRICT OUTPUT FORMAT Return only a JSON array of your answer. Each array element must be an object with keys ‘"id"‘, ‘"issue"‘ and ‘"evidence"‘. Example: ‘‘‘json [ {{"id": 1,"issue": "...", "evidence": "..."}}, {{"id": 2,"issue": "...", "evidence": "..."}} ] ‘‘‘ Guidelines: - Make sure to identify distinct and independent issues (e.g., missing constraints, wrong variable bounds, or incorrect objective formulation). - Do NOT include issues that do not directly affect the models ability to reach optimality. - Only output the JSON array. DO NOT include any explanations, markdown, or extra text before or after the JSON array. Now take a deep breath and think step by step.
Diagnose positive and negative tasks of an insight You are an expert in Industrial Engineering and Operations Research. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague which failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as *the failed mathematical model*) 3. A diagnostic report on the proposed mathematical model, identifying all formulation issues that prevent optimality. 4. A collection of insights your colleague previously consulted to generate the mathematical model, each insight includes: - insight_id: the unique ID for this insight - condition: the problem-specific context and broader modeling situations in which the insight should apply to avoid mistakes (i.e., its applicability condition) - explanation: the description of the pitfalls and guidance for the proper principle/practice - example: the demonstration showing wrong vs. correct version (principle, formula, or code snippet) ### Problem description {problem_description} ### The failed mathematical model Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘), with brief comments indicating the corresponding insight_id and how it helps the specific formulation. {failed_formulation} ### The diagnosed issues {diagnosed_issues} ### A collection of insights {retrieved_insights} ### Your task Step 1: Carefully review the problem description, and analyse the failed mathematical model with the issues identified in the diagnostic report. Step 2: Examine the collection of insights and the corresponding annotations in the proposed model that indicate how each insight was implemented. Step 3: For each insight, determine its correctness by comparing with the gold-standard program, then evaluate its implementation impact (whether it leads to any of the identified issues): 1. Assign "positive" label when: The insight is correct and adopted by the model, and it contributed to a correct formulation (helped achieve the right result). 2. Assign "invalid" label when: The insight is correct but not adopted; if adopted, it would have helped produce the correct formulation and resolve identified issues. 3. Assign "negative" label when: - The insight is wrong and adopted, thereby directly causing an identified issue; - The insight is wrong and not adopted, yet it provides inapplicable/misleading guidance that poses a risk of errors. 4. Assign "irrelevant" label when: The insight is irrelevant to the mathematical modeling in this optimization task and did not affect your colleagues formulation. Suggested decision order: - Judge correct vs. wrong. - Check adopted vs. not adopted. - Assess impact (enabled correctness, caused issues, could resolve issues, risk/irrelevant). - Map to one of the four labels. Step 4: Record the insight_id and the assigned label. Cite concrete evidence from the problem description and diagnosed issues, and justify the labeling. Clearly explain the mapping between each insight and the formulation issues. ### STRICT OUTPUT FORMAT Return only a JSON array of your answer in Step 4. Each array element must be an object with keys ‘"insight_id"‘ (integer), ‘"state"‘("postive" or "negative") and ‘"evidence"‘ (string). Example: ‘‘‘json [ {{"insight_id": 1, "state": "positive", "evidence": "<text>"}}, {{"insight_id": 5, "state": "negative", "evidence": "<text"}} ] ‘‘‘ Guidelines: - Make sure to identify and output each insight_id with its state. Do NOT miss any insight id in the given collection. - Only output the JSON array. DO NOT include any explanations, markdown, or extra text before or after the JSON array. Now take a deep breath and think step by step.
Diagnose unretrieved tasks of an insight You are an expert in Industrial Engineering and Operations Research. You are given: 1. A problem description for the optimization task 2. A mathematical model proposed by your colleague which failed to yield an optimal solution when solved with the Gurobi optimizer (hereafter referred to as *the failed mathematical model*) 3. A diagnostic report on the proposed mathematical model, identifying the specific issue that prevents optimality 4. A collection of insights. Each insight includes: - insight id: the unique identifier of the insight - taxonomy: the classification of the modeling/formulation/code-implementation aspect it pertains to - condition: the problem-specific context and broader modeling situations in which the insight applies (its applicability condition) ### Problem description {problem_description} ### The failed mathematical model Note: the model is written in LaTeX and presented in a plain-text code block (‘‘‘) {failed_formulation} ### The diagnosed issue {one_diagnosed_issue} ### Candidate insights {candidate_insights} ### Your task Step 1: Carefully review the problem description, and analyse the failed mathematical model with the issue identified in the diagnostic report. Step 2: Evaluate each candidate insight individually. Retain only those that directly apply to resolving the identified issue in the diagnostic report. For every insight, cite concrete evidence from the problem description and diagnosed issues, and justify how the insight helps fix the specific modeling issue. Step 3: Rank the applicability of the selected insights in descending order. ### STRICT OUTPUT FORMAT Return only a JSON array of your result from Step 3. Each array element must be an object with keys ‘"insight id"‘ (integer), ‘"ranking"‘(applicability rank; 1 = highest) and ‘"evidence"‘ (string). Example: ‘‘‘json [ {{"insight_id": 1, "ranking": 1, "evidence": "<text>"}}, {{"insight_id": 5, "ranking": 2, "evidence": "<text>"}}, {{"insight_id": 7, "ranking": 3, "evidence": "<text>"}} ] ‘‘‘ Guidelines: - Output only the insights that apply to the identified issue(s). - Only output the JSON array. DO NOT include any explanations, markdown, or extra text before or after the JSON array. Now take a deep breath and think step by step. You will be awarded a million dollars if you get this right.
Refine Insight Conditions You are an expert in Industrial Engineering and Operations Research. Your task is to design multiple refinement strategies for the condition of a given insight to improve its applicability in optimization tasks. You are given: 1. Original insight with three fields: - condition: Trigger specifying when the insight applies, grounded in problem description/domain features. It first states the general situation, then illustrates with the specific problem. - explanation: Under this condition, the description outlines the best practice, the common mistake and its cause. It illustrates the issue with this problem as an example and generalizes the correct modeling strategy it reflects. - example: Wrong vs. correct demonstration (principle, formula, or code). 2. Task-derived insight conditions: - Inapplicability conditions of insights for negative tasks where prior use of the insight misled the modeling. - Applicability conditions of insights for unretrieved tasks which should have retrieved these insights but were missed. Applicability conditions for unretrieved tasks where the insight was relevant but not retrieved. ### The original insight to be refined {original_insight} ### inapplicability conditions from negative tasks {neg_conditions} ### applicability conditions from unretrieved tasks {unr_conditions} ### Your tasks Your goal is to refine only the condition field of the original insight so that it excludes as many negative tasks as possible, captures as many previously unretrieved tasks as possible, and still applies to the tasks covered by the original condition. To achieve this, consider the following four steps: Step 1: Consolidate inapplicability: Merge all inapplicability conditions from negative tasks into a single, unified inapplicability condition. Use the pattern: "This insight does NOT apply when [general situation that negates the insight]. For example, when the problem statement mentions [concrete trigger(s) grounded in the problem description or defining features indicating properties that conflict with the insight]." Step 2: Consolidate applicability: Merge all applicability conditions from unretrieved tasks into a single, unified applicability condition. Use the pattern: "This insight applies when [general situation that warrants the insight]. For example, when the problem statement mentions [concrete trigger(s) grounded in the problem description or defining features indicating properties that align with this insight]." Step 3: Integrate into the insights condition field: - Preserve the original condition text verbatim. - Structure the field as three paragraphs using the pattern below. - Use the unified inapplicability condition as the second paragraph. Output using the pattern (three brief paragraphs): - First paragraph: original applicability condition - Second paragraph: unified applicability condition for unretrieved tasks: "This insight applies when ..." - Third paragraph: unified inapplicability condition for negative tasks: "This insight does NOT apply when ..." Step 4: Generate {path_k} distinct refinement strategies (paths): Generate distinct refinement strategies (paths) for how you consolidate the insight conditions. For each path, write applicability/inapplicability using the required pattern in Step 3. Examples: - one broad rule: Merge multiple situations into a single general trigger that clearly and broadly covers the main applicable scenario. - short OR list: Provide a brief list of alternative triggers; if any one appears, the insight applies. - must-have pair: Require two key cues to occur together before the insight applies, reducing false positives. - info stated vs. missing: Apply only when the problem explicitly states the required details; treat omissions as not applicable. - helpful keywords: Anchor applicability on a small set of representative keywords or phrases that reliably signal the scenario. - narrow the scope: Add concise qualifiers to limit coverage to a well-bounded context so the insight applies only under those limits. ### STRICT OUTPUT FORMAT Only Return a JSON array with the following structure: [ { "path_id": 1, "strategy": "<1-2 sentences description of the refinement strategy>", "new_condition": "<rewritten condition string>" }, { "path_id": 2, "strategy": "...", "new_condition": "..." }, ... ] ### Guidelines - Always output valid JSON. - Do not include explanations outside the JSON. - Each path must be semantically distinct. Now take a deep breath and think step by step.

### v3

@@
-AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library
-Abstract
-Optimization modeling enables critical decisions across industries but remains hard to automate: informal language must be mapped to precise mathematical formulations and executable solver code, while prior LLM approaches either rely on brittle prompting or costly retraining with limited generalization.
-We present AlphaOPT, a self-improving experience library that enables an LLM to learn from limited demonstrations (i.e, even answers alone without gold-standard program) and solver feedback without annotated reasoning traces or parameter updates.
-AlphaOPT operates a continual two-phase cycle: (i) a Library Learning phase that reflects on failed attempts, extracts solver-verified, structured insights as $\{\textit{taxonomy},\ \textit{condition},\ \textit{explanation},\ \textit{example}\}$; and (ii) a Library Evolution phase that diagnoses retrieval misalignments and refines the applicability conditions of stored insights, improving transfer across tasks.
-This design (1) learns efficiently from limited demonstrations without curated rationales, (2) expands continually without costly retraining by updating the library rather than model weights, and (3) makes knowledge explicit and interpretable for human inspection and intervention.
-Experiments show that AlphaOPT steadily improves with more data (65% $\rightarrow$ 72% from 100 to 300 training items) and surpasses the strongest baseline by 7.7% on the out-of-distribution OptiBench dataset when trained only on answers.
+AlphaOPT: Formulating Optimization Programs with a Self-Improving LLM Experience Library
+Abstract.
+Optimization modeling underlies critical decision-making across industries, yet remains difficult to automate: natural-language problem descriptions must be translated into precise mathematical formulations and executable solver code.
+Existing LLM-based approaches typically rely on brittle prompting or costly retraining, both of which offer limited generalization.
+Recent work suggests that large models can improve via experience reuse, but how to systematically acquire, refine, and reuse such experience in structurally constrained settings remains unclear.
+We present AlphaOPT, a self-improving experience library that enables LLMs to learn optimization modeling knowledge from limited supervision, including answer-only feedback without gold-standard programs, annotated reasoning traces, or parameter updates.
+AlphaOPT operates in a continual two-phase cycle: a Library Learning phase that extracts solver-verified, structured insights from failed attempts, and a Library Evolution phase that refines the applicability of stored insights based on aggregate evidence across tasks.
+This design allows the model to accumulate reusable modeling principles, improve transfer across problem instances, and maintain bounded library growth over time.
+Evaluated on multiple optimization benchmarks, AlphaOPT steadily improves as more training data become available (65% $\rightarrow$ 72% from 100 to 300 training items) and outperforms the strongest baseline by 9.1% and 8.2% on two out-of-distribution datasets.
+These results demonstrate that structured experience learning, grounded in solver feedback, provides a practical alternative to retraining for complex reasoning tasks requiring precise formulation and execution.
 AlphaOPT code and data are available at https://github.com/Minw913/AlphaOPT.
-1 INTRODUCTION
-Optimization models support critical decision-making in finance, manufacturing, marketing, transportation, and logistics (AhmadiTeshnizi et al., 2023; Bertsimas & Tsitsiklis, 1997; Ramamonjison et al., 2022).
-Beyond improving efficiency, automating the optimization workflow lowers the barrier to operations research expertise in industry, enabling non-experts to prototype faster, iterate on formulations, and deploy solver-backed decisions at scale.
-Yet this process has long been challenging, as informal and often ambiguous specifications must be mapped to precise, domain-specific formulations and paired with appropriate code and solvers, creating major bottlenecks for end-to-end automation (Jiang et al., 2025).
+Large Language Models, Optimization Problems, Experience Learning
+$c$s: Applied computing Operations researchccs: Computing methodologies Artificial intelligence
+1. Introduction
+Optimization models support critical decision-making in finance, manufacturing, marketing, transportation, and logistics (AhmadiTeshnizi et al., 2024; Bertsimas and Tsitsiklis, 1997; Ramamonjison et al., 2022).
+Beyond improving efficiency, automating the optimization workflow lowers the barrier to operations research expertise in industry and enables non-experts to prototype faster, iterate on formulations, and deploy solver-backed decisions at scale.
+Yet this process has long been challenging, because informal and often ambiguous specifications must be mapped to precise, domain-specific formulations and paired with appropriate code and solvers.
+This creates major bottlenecks for end-to-end automation (Jiang et al., 2025).
 Advances in large language models (LLMs) make this vision increasingly feasible: They can parse natural language requirements (Ouyang et al., 2022), generate executable programs (Nijkamp et al., ; Jimenez et al., 2024), and orchestrate downstream tools (Qin et al., 2024).
 Two main lines of work have emerged.
-Prompt-based systems steer general LLMs with structured prompts and tool use (Xiao et al., 2023; Thind et al., 2025; AhmadiTeshnizi et al., 2024; Zhang & Luo, 2025).
+Prompt-based systems steer general LLMs with structured prompts and tool use (Xiao et al., 2023; Thind et al., 2025; AhmadiTeshnizi et al., 2024; Zhang and Luo, 2025).
 Fine-tuning approaches adapt models on domain corpora and benchmarks (Huang et al., 2025; Yang et al., 2024).
-Despite this progress, both families face limitations: prompt-based systems stop improving once they run out of fixed templates, and they are fragile to small wording changes and shifts in the domain; fine-tuned models require costly retraining and, critically, most benchmarks and datasets in the community (e.g., NLP4LP (AhmadiTeshnizi et al., 2024), MAMO (Huang et al., 2024), IndustryOR (Huang et al., 2025)) contain only programs/solutions rather than the intermediate reasoning that governs modeling choices, thereby limiting the generalizability of fine-tuning approaches.
+Despite this progress, both families face important limitations.
+Prompt-based systems improve only within the scope of their fixed prompt templates and are sensitive to small wording changes and domain shifts.
+Fine-tuned models, on the other hand, require retraining as new data arrives and, critically, most benchmarks and datasets in the community (e.g., NLP4LP (AhmadiTeshnizi et al., 2024), MAMO (Huang et al., 2024), IndustryOR (Huang et al., 2025)) provide only final programs or solutions rather than the intermediate reasoning that governs modeling decisions, which limits the generalizability of fine-tuning approaches.
 This motivates a new learning paradigm for optimization formulation: Instead of relying solely on prompts or retraining, LLMs should continually improve by accumulating, refining, and reusing solver-verified modeling insights.
-Figure 1: AlphaOPT identifies and retrieves relevant insights to guide problem solving.
-In this example, it avoids the common mistake of minimizing the sum of process times and instead introduces an auxiliary variable to correctly minimize the makespan, leading to the correct solution.
-We propose AlphaOPT, a self-improving framework that builds and refines a structured library of solver-verified insights for optimization formulation, as exemplified in Figure 1.
-Each insight encodes a reusable modeling rule in the form of a 4-tuple (taxonomy, (applicability) condition, explanation, example), which specifies not only what to reuse but also when and why it applies.
-We remark that our library learning framework does not require backpropagation to update framework parameters and can be regarded as the evolutionary mechanism.
-AlphaOPT improves through a continual two-phase cycle.
-Library Learning acquires new insights from both gold programs (when available) and solver-verified answer-only supervision, organizing them into a dynamically updating hierarchical taxonomy.
-Library Evolution then diagnoses misalignments between tasks and insight applicability, and refines conditions using aggregate evidence, ensuring that insights remain neither too narrow nor overly general.
-This design yields a principled optimization view: library construction corresponds to maximizing expected task success induced by task–insight matching while regularizing size to maintain efficiency and prevent redundancy.
-We conduct quantitative experiments across multiple benchmarks and baselines, as well as qualitative analyses of the learned library.
-The results show that, compared to prompt-based or fine-tuning approaches, AlphaOPT (1) learns efficiently from limited demonstrations (i.e., it can learn from answers without recalling formulation) without requiring annotated reasoning traces or even gold-standard programs, (2) achieves stronger out-of-distribution generalizability and more consistent continual growth than learning-based methods, and (3) makes knowledge explicit and interpretable for human inspection and involvement.
-Beyond these advantages, AlphaOPT also achieves state-of-the-art performance on multiple benchmarks.
-These results demonstrate the efficacy and potential of self-improving experience-library learning for optimization formulation, paving the way toward more challenging settings, such as efficient program formulation and large-scale optimization.
-Our main contributions are threefold:
-  • A library learning framework that learns from answers only.
-We propose the first experience-library learning framework for natural language optimization formulation tasks, formally grounded in a mathematical view.
-The system can learn solely from answers, without requiring gold-standard programs.
-  • A reusable and interpretable experience library.
-We construct the first solver-verified library of structured modeling insights for LLM systems, designed to be reusable across tasks and explicitly interpretable for reliable transfer in operations research domains.
-  • State-of-the-art out-of-distribution generalization.
-AlphaOPT achieves strong generalization beyond training distributions, attaining state-of-the-art performance on LogiOR and OptiBench benchmarks.
-2 Related Work
-LLMs for Solving Optimization Problems.
-Related work can be categorized into prompt-based and learning-based approaches.
-Prompt-based methods guide reasoning and modeling through multi-step prompts using proprietary LLMs (AhmadiTeshnizi et al., 2024; Xiao et al., 2023).
-AhmadiTeshnizi et al.
-(2023) first introduced OptiMUS, demonstrating how LLMs can generate optimization formulations from natural language descriptions, and OptiMUS-0.3 (AhmadiTeshnizi et al., 2024) extends this line of work to large-scale instances, introducing retrieval-augmented prompting and solver-integrated verification to improve scalability.
-In contrast, learning-based methods construct training datasets and apply instruction tuning to open-source LLMs.
-Huang et al.
-(2025) proposed a semi-automated pipeline to synthesize training data, which is then used to fine-tune an open-source ORLM model.
-LLMOPT (Jiang et al., 2025) combines both paradigms by modeling optimization with five elements and fine-tuning on expert-annotated data via multi-instruction SFT.
-More recently, ORThought (Yang et al., 2025a) introduced the LogiOR benchmark and an expert-guided chain-of-thought framework, providing a systematic dataset and evaluation pipeline for optimization tasks that require harder logic.
-In terms of multi-agent design, Xiao et al.
-(2023) employs a collaborative multi-expert framework to enhance reasoning, Zhang & Luo (2025) integrates sandbox-based code execution and self-repair/self-verification.
-Several benchmarks now exist that cover LP, MILP, NLP, and other optimization problem types (Xiao et al., 2023; AhmadiTeshnizi et al., 2024; Huang et al., 2025; Yang et al., 2024).
-Yet, none of the prior work has investigated strengthening LLMs’ optimization capabilities by learning and reusing structured modeling experience.
-Decision-making tasks with Library Learning.
-Library Learning refers to the process where reusable patterns or modules are automatically extracted from past experiences to improve future problem-solving.
-These experiences include concrete trajectories or demonstrations, as well as abstracted rules generalized from successful or failed attempts (Zhao et al., 2024; Mu et al., 2025; Feng et al., 2025; Wang et al., ; Zhu et al., 2023).
-In terms of experience improvement, Zhao et al.
-(2024) and Mu et al.
-(2025) leverage an LLM to prune the library by checking if a newly added insight duplicates or conflicts with existing insights, or merges and generalizes from those overlapping insights.
-Feng et al.
-(2025) uses check functions to ensure that LLM-translated action sequences remain within the generalization boundaries of the original experiences.
-LLM-driven Evolutionary Methods.
-Recent LLM-driven evolutionary frameworks have achieved remarkable advances in scientific discovery, showcasing LLM’s capacity for broad generative exploration on solutions.
-Romera-Paredes et al.
-(2024) introduces FunSearch, a genetic programming driven by LLMs to search for feasible or optimal solutions of mathematical problems.
-AlphaEvolve (Novikov et al., 2025) extends the FunSearch system to provide the ability to perform multiobjective optimization using rich forms of natural-language context and feedback.
-Grayeli et al.
-(2024) applies LLMs to discover abstract concepts from high-performing hypotheses, combining symbolic regression with LLM-guided exploration within a concept library.
-ReEvo (Ye et al., 2024) frames LLMs as hyper-heuristics with a reflective evolution mechanism, enabling the generation and iterative refinement of heuristics across multiple NP-hard problems.
-HeurAgenix (Yang et al., 2025b) further develops this direction by evolving a pool of heuristics and dynamically selecting the most suitable one for each problem state, highlighting LLMs’ role in adaptive heuristic discovery.
-Besides, LLM-guided evolution has also found use in discovering heuristics for combinatorial optimization (Liu et al., ).
-3 Methodology
-Optimization tasks arrive with diverse natural-language descriptions, yet they share recurring modeling rules that activate under identifiable conditions.
-We identify three major challenges in building reliable systems that leverage LLMs to formulate and solve optimization problems using existing technologies and resources.
-First, gold-standard programs are scarce and may contain annotation errors (Jiang et al., 2025; Yang et al., 2025a), while datasets with only answer labels remain underutilized (Huang et al., 2024; 2025; Lu et al., 2025).
-Second, fine-tuned models (Huang et al., 2025; Jiang et al., 2025) struggle to generalize because the crucial when-to-apply-what knowledge is weakly represented (or missing) in training data; they can mimic syntax without mastering applicability.
-Third, the performance of prompt-based agent systems AhmadiTeshnizi et al.
-(2023); Xiao et al.
-(2023); Yang et al.
-(2025a) stagnates as the number of exemplars increases: they rely on human empirical curation and lack the capacity to adapt or to continually learn from larger datasets.
-We propose AlphaOPT, an experience-library learning framework that learns from both gold programs (when available) and answer-only supervision.
-AlphaOPT iteratively builds a structured, solver-verified repository of reusable insights with explicit applicability conditions and evolves these conditions at the population level to improve generalization while avoiding redundancy.
-This two-stage design is described in Section 3.1.
+Indeed, recent approaches such as Reflexion (Shinn et al., 2023), STaR (Zelikman et al., 2022), ExpeL (Zhao et al., 2024), and AlphaEvolve (Novikov et al., 2025) demonstrate that large models can improve through experiential reuse, storing reflections, rationales, or code edits and applying them in new tasks.
+These methods have been effective in common reasoning tasks such as information retrieval (Yang et al., 2018) and web browsing (Yao et al., 2022), but they face limitations for optimization problems.
+First, their experiences are largely preserved as free-form text or edits, without explicit applicability semantics or refinement mechanisms to represent, verify, and refine applicability-sensitive modeling knowledge; in optimization tasks, applying mismatched experiences can therefore have detrimental effects.
+Second, when generating experiences, these methods often lack rigorous replay-based validation, or rely on verifications limited to task-level outcomes (e.g., rewards, final answers, or test-case success), which does not guarantee that the underlying knowledge is structurally valid or transferable.
+Third, their experiences tend to be organized as a flat list of items, whereas operations research involves diverse insights across different modeling stages and problem types, inherently requiring a more efficient organization of acquired knowledge.
+Regarding these challenges, we introduce AlphaOPT, an experience learning framework for optimization formulation that builds and continually refines a library of solver-verified modeling insights.
+In contrast to prior approaches that store experience as unstructured text or static edits, AlphaOPT formulates experience as structured, reusable insights with explicit applicability semantics.
+Each insight is defined as a four-tuple(taxonomy, condition(applicability), explanation, example), specifying not only what modeling rule to reuse, but when and why it should be applied given a natural-language problem description.
+AlphaOPT improves performance through a continual two-phase cycle without updating model parameters.
+In the Library Learning phase, the framework self-explores optimization tasks and autonomously constructs a hierarchical taxonomy to organize insights across modeling stages, while relying on solver verification to ensure correctness even in the absence of gold-standard programs.
+In the Library Evolution phase, AlphaOPT diagnoses systematic misalignments between tasks and insights and refines applicability conditions based on aggregate evidence, preventing insights from being either overly narrow or overly general.
+Compared to existing experience learning methods, AlphaOPT introduces three key advances: (1) a self-updating hierarchical taxonomy for organizing insights; (2) solver-verified insight generation and updating; and (3) an aggregate-driven refinement mechanism that continuously improves the precision and generalizability of insight applicability.
+We conduct quantitative experiments across multiple benchmarks and baselines.
+Compared with other approaches, the results show that AlphaOPT not only achieves state-of-the-art performance on multiple benchmark datasets without any gold-standard formulation or program, but also demonstrates strong generalization on out-of-distribution datasets, and continual performance growth with converging library size.
+These results demonstrate the efficacy of experience learning for optimization formulation, which paves the way toward more challenging settings, such as efficient program formulation and large-scale optimization.
+Compared to existing approaches, AlphaOPT makes the following contributions:
+  • We introduce the first experience learning framework for natural-language optimization modeling, which represents solver-verified modeling knowledge as condition-aware insights that are continually refined over time.
+  • Our experiments show that AlphaOPT generalizes strongly beyond training distributions, achieving state-of-the-art performance on the LogiOR and OptiBench benchmarks.
+  • We provide extensive qualitative and quantitative analyses of the learned experience library and its refinement dynamics, demonstrating explicit, interpretable, and human-inspectable knowledge transfer in operations research domains.
+Figure 1.
+AlphaOPT iteratively learns an experience library through two phases: a Library Learning phase that extracts structured insights from failed attempts and organizes them in a hierarchical taxonomy, and a Library Evolution phase that refines insight applicability based on all associated tasks avoid over- or under-generalization.
+The figure shows a two-phase process for learning and refining optimization knowledge.
+On the left, Phase 1 illustrates library learning.
+The process begins with an optimization task, including a natural language problem statement, a known correct answer, and a gold-standard mathematical program.
+The system attempts to solve the task, learns from both mistakes and successful solutions, and uses these attempts to extract reusable experience.
+This experience is organized into a structured taxonomy that covers domain modeling, general mathematical formulation, and code implementation.
+From this structure, the system generates concrete insights, each with a clear condition describing when it applies, an explanation of the modeling principle, and examples that contrast incorrect and correct formulations.
+On the right, Phase 2 illustrates library evolution.
+Previously generated insights are examined by comparing cases where they apply and where they do not.
+Using performance and error signals aggregated across tasks, the system refines the conditions under which each insight should be used.
+The result is a more precise description of applicability, for example distinguishing problems that require closed vehicle tours with flow balance from problems that allow open or multi-day routes.
+Overall, the figure shows how raw optimization attempts are transformed into structured, conditional knowledge and how that knowledge is iteratively refined to improve preciseness of the applicability and transfer across diverse problems.
+2. Related Work
+2.1. LLMs for Solving Optimization Problems
+Recent work increasingly uses LLMs as autonomous agents for optimization, mainly following two paradigms: formulation and heuristic generation (Ye et al., 2024; Yang et al., 2025b; Wu et al., 2025).
+LLM-driven formulation tasks aim to translate natural-language problem statements into solver-ready mathematical programs and executable solving code.
+Multiple benchmark datasets have been proposed to cover diverse problem families such as LP, MILP, and NLP (Xiao et al., 2023; AhmadiTeshnizi et al., 2024; Yang et al., 2024, 2025a).
+Existing approaches can be broadly grouped into prompt-based and training-based methods.
+Prompt-based methods use multi-step prompting and structured reasoning to guide LLMs in parsing problems, extracting variables/constraints, and synthesizing models.
+Representative examples include Chain-of-Experts (Xiao et al., 2023), which leverages multi-role decomposition, and OptiMUS (AhmadiTeshnizi et al., 2024), which employs a self-reflect agentic pipeline with retrieval-augmented prompting to produce mathematical formulations.
+LEAN-LLM-OPT (Liang et al., 2026) further integrates a three-agent workflow with RAG and few-shot learning for problem typing, example retrieval, and data grounding.
+However, these methods lack sustained knowledge injection and long-term memory, which can lead to performance plateaus under distribution shift.
+Training-based methods improve robustness by building instruction corpora and fine-tuning open-source models.
+For example, ORLM (Huang et al., 2025) synthesizes natural-language optimization problem data used for fine-tuning, and LLMOPT (Jiang et al., 2024) adopts a five-element schema with multi-instruction SFT on expert-labeled and augmented data.
+Recent post-training work further applies RL-style refinement, such as OR-R1 (Ding et al., 2025), to improve single-shot performance and robustness.
+Nonetheless, these pipelines rely heavily on high-quality, diverse data, and are often hard to interpret or audit.
+Overall, despite progress in prompting, data synthesis, fine-tuning, and post-training, existing methods remain insufficient for reliable deployment and sustained improvement in real-world operations research systems.
+2.2. Decision-making Tasks with Experience Learning
+Experience Learning studies how LLM-based agents extract and reuse transferable structured experience from historical interactions to continuously improve downstream decision-making and execution (Zhao et al., 2023; Shinn et al., 2023).
+Such experience can take the form of episodic traces (Park et al., 2023) (e.g., tool/action sequences), semantic insights abstracted from successful or failed attempts (Zhao et al., 2023).
+This line of work is primarily evaluated on long-horizon, interactive, feedback-rich decision-making tasks, including multi-tool planning and web interaction (Qin et al., 2023; Zhou et al., 2023), repository-level program repair/code generation (Jimenez et al., 2023), and multi-step reasoning problem solving (Shinn et al., 2023; Wang et al., 2024).
+Shinn et al.
+(2023) propose Reflexion, which enables language agents to learn from trial-and-error by converting task feedback into verbal self-reflections that are stored as episodic memory and reused in subsequent attempts.
+ExpeL (Zhao et al., 2023) jointly extracts similar exemplars/trajectories together with their associated insights to guide task execution and further employs a voting-based mechanism to downweight mutually contradictory insights, thereby improving single-shot performance and cross-task generalization.
+At a more symbolic end, Zhu et al.
+(2023) explicitly induces a rule library and maintains it using quantitative filters such as coverage and confidence, enhancing controllability and robustness of experience reuse.
+AgentKB (Tang et al., 2025) and ReasoningBank (Ouyang et al., 2025) move toward organizing an evolving reasoning memory to support self-improving agents across tasks.
+However, current frameworks face systematic limitations in three key aspects.
+First, regarding experience applicability, most methods still rely on embedding-based similarity retrieval or coarse metadata to match experiences (Zhao et al., 2023; Park et al., 2023; Wang et al., 2024), while lacking explicit, verifiable preconditions (when an experience applies, when it should not, and what assumptions it requires), which increases the risk of negative transfer under distribution shift.
+Second, for strict verification of experience generation, merging, and editing, existing mechanisms often depend on LLM self-assessment (Zhao et al., 2023; Shinn et al., 2023; Feng et al., 2025); for instance, ExpeL’s voting-based filtering can remove contradictions but does not guarantee semantic correctness.
+Finally, regarding cross-task generalization, existing experience libraries typically support experience accumulation and local pruning (Zhao et al., 2023; Tang et al., 2025; Ouyang et al., 2025), but lack a systematic mechanism to assess whether an experience that helps on local source tasks may degrade performance on broader related task sets.
+3. Methodology
+Building reliable LLM-based systems for solving optimization problems remains challenging.
+To begin with, gold-standard formulations and programs are scarce and may contain annotation errors (Jiang et al., 2025; Yang et al., 2025a), while datasets with only answer-level supervision are underexploited (Huang et al., 2024, 2025; Lu et al., 2025).
+Moreover, although optimization tasks arrive with diverse natural-language descriptions, they often share recurring modeling rules that activate under identifiable conditions; however, existing methods still struggle to learn such when-to-apply-what knowledge.
+Fine-tuned models (Huang et al., 2025; Jiang et al., 2025) often overfit to syntax without mastering applicability, whereas prompt-based agent systems (AhmadiTeshnizi et al., 2024; Xiao et al., 2023; Yang et al., 2025a) rely on human empirical curation and lack the capacity to adapt or continually learn from larger datasets.
+To this end, we propose AlphaOPT, which constructs an experience library of reusable insights with explicit applicability conditions, and then iteratively evolves these conditions at population level to improve generalization.
+As illustrated in Figure 1, it consists of two complementary phases that form a continual cycle of acquisition and refinement.
+The first phase, Library Learning, extracts structured insights from individual failed tasks under answer-only supervision while minimizing redundancy.
+The second phase, Library Evolution, diagnoses misalignments between insights and tasks and refines applicability conditions to enhance generalization while reducing confusion caused by overgeneralization.
 In Section 3.3, we provide a mathematical interpretation that frames library construction as maximizing task success with a size regularizer.
-In Section A, we compare our method with prior works on learning from experience and self-evolving problem-solving agents.
-3.1 AlphaOPT Framework
-Figure 2: AlphaOPT builds and refines an experience library over multiple iterations.
-Each iteration consists of a library learning stage, which produces insights from attempts to solve tasks, and a library refinement stage, which adjusts the applicability of insights to avoid being too narrow or too general.
-The framework incrementally learns a structured library of experiences over iterations until a stopping criterion indicates that the current model can no longer make meaningful improvements.
-As illustrated in Figure 2, each iteration consists of two complementary phases that form a continual cycle of acquisition and refinement.
-The first phase, Library Learning, extracts insights from individual tasks under either gold-program or answer-only supervision while minimizing redundancy.
-The second phase, Library Evolution, diagnoses misalignments between insights and tasks and refines applicability conditions to enhance generalization while reducing confusion caused by overgeneralization.
-The design follows three guiding principles: it is failure-driven (every error becomes a learning opportunity), locally validated (an insight must solve its source task before being admitted), and compact yet generalizable (redundant insights are merged and conditions refined to prevent uncontrolled growth that hinders retrieval and execution).
-The prompts for all LLM modules are provided in Appendix E.
-Figure 3: The overall workflow of library learning and evolution.
-The left panel depicts two complementary flows: library learning, which extracts new insights by generating and consolidating insights from failed optimization attempts, and library diagnosis, which analyzes interactions between failed tasks and retrieved insights to collect negative and unretrieved cases for refinement.
-The right panel illustrates library refinement, where the LLM refines each insight’s applicability conditions, validates them via retrieval replay, and integrates the updated insights back into the library.
-3.1.1 Library Learning
-The objective of this stage is to generate reusable insights as structured 4-tuples (Taxonomy, Condition, Explanation, Example) and organize them in a hierarchical taxonomy for efficient retrieval, while minimizing redundancy in the library.
-The left panel of Figure 3 (see the Library learning flow label) illustrates the workflow for this stage.
-Insight Extraction, Representation, and Supervision Mode.
-Insights can be learned from either problems with a gold-standard program or with the answer alone.
-For each task, the system first constructs a mathematical formulation, then generates an executable solver program, and invokes the solver.
+Implementation details for the different phases are provided in the Appendix A.
+To the best of our knowledge, our framework is the first to adapt experience learning to operations research (OR) with three key innovations: (1) structured knowledge for interpretability and auditability: Each insight is represented with taxonomy, condition, explanation, and example, which renders its applicability explicit, reviewable, and even revisable in practice; (2) solver-guided verifiability: If a program generated by LLM achieves the optimal objective under the solver, it is highly likely to be valid and can serve as a reliable anchor for extracting insights, which broadens the sources of experience collection.
+Newly generated insights are retest to be retrievable in the library and to succeed when replayed on its source task, which ensures that they are valid before integration; (3) refinement of experience applicability for generalizability and preciseness: Applicability conditions are refined using cross-task evidence, so insights neither over-generalize nor remain too narrow, which improves safe transfer across problem families.
+Figure 2.
+The overall workflow of AlphaOPT.
+The left panel depicts two complementary flows: library learning, which self-explores correct patterns and extracts new insights by trial and error; library diagnosis, which categorizes interactions between failed tasks and retrieved insights into positive, negative, and unretrieved cases.
+The right panel shows library refinement, whereby the LLM refines insight applicability, validates via retrieval replay, and reintegrates them into the library.
+The figure illustrates how the insight library is learned, diagnosed, and refined through interaction with optimization tasks.
+On the left, the library learning and diagnosis process is shown.
+Given an optimization task, the system retrieves relevant prior insights based on a taxonomy and their applicability conditions.
+These insights guide the construction of a mathematical formulation, which is then translated into a solving program.
+If the program succeeds, the task provides positive evidence for the retrieved insights.
+If it fails, the system performs self-exploration by comparing the failed solution with a gold-standard program and extracting new insights.
+All insights, both existing and newly generated, are stored in the library.
+At the same time, the system diagnoses insight applicability across related tasks.
+For each task, an insight can be observed as helpful, harmful, or not retrieved at all.
+Insights that consistently perform poorly or are missing when needed are marked as candidates for refinement.
+On the right, the library refinement process is shown.
+Candidate insights are replayed through library retrieval and self-verification on source tasks.
+Based on positive, negative, or unretrieved outcomes, their applicability conditions are updated by keeping, removing, or adding conditions.
+The refined versions are validated, and the best-performing versions are selected to form the updated library.
+Overall, the figure shows a closed loop in which solving optimization tasks both uses the library and continuously improves it by refining when each insight should apply.
+3.1. Library Learning
+This stage is designed to generate reusable insights as structured 4-tuples (Taxonomy, Condition, Explanation, Example) and organize them in a hierarchical taxonomy dictionary for efficient retrieval while minimizing redundancy in the library.
+Figure 2 (Library Learning Flow) clearly illustrates the overall workflow.
+Insight Extraction and Representation
+For each arriving task, the system first constructs a mathematical formulation, then generates an executable solver program and invokes the solver.
 When the library is non-empty, both steps are guided by retrieved insights.
-If the generated program does not achieve the correct optimal value, two supervision modes are used to guide the generation of new insights.
-When a gold program is available, the system compares the candidate formulation and program against the reference, diagnosing discrepancies (e.g., missing variables, misformulated constraints, incorrect objective terms) and distilling them into insights.
-When only the answer (i.e., final optimal objective) is provided, the system performs solver-guided self-exploration: it iteratively proposes executable programs, reuses prior failures as context, and receives verification from the solver.
+If the generated program does not achieve the correct optimal value, the system performs solver-guided self-exploration: It iteratively proposes executable programs, reuses prior failures as context, and receives verification from the solver.
 Once a program achieves its correct objective, it is treated as a proxy for the gold standard in anchor insight extraction.
-Before being stored in the library, each insight is locally verified by reapplying it to its source task to ensure that it resolves the original failure.
-In addition, to mitigate stochastic successes that could obscure useful lessons, we conduct three independent trials per task, allowing errors from probabilistic generation to serve as signals for learning.
-Each insight is represented as a structured 4-tuple: Taxonomy, hierarchical labels for indexing and retrieval; Condition, an explicit description of the applicability signals in the problem; Explanation, the underlying principle of applying this insight; and Example, a concrete demonstration such as a mathematical constraint or code snippet.
-Library Storage and Retrieval.
-Insights are stored in a dynamically updated hierarchical taxonomy organized into three main tracks: Domain Modeling (problem-specific structures and assumptions), General Formulation (reusable mathematical patterns), and Code Implementation (solver-specific coding practices).
-Under each track, insights are further classified with two-level labels, where Level-1 captures a broad category and Level-2 refines it into a more specific subcategory.
-The taxonomy is initialized with few-shot labels and expands online: each new insight is either mapped to an existing category or, if no suitable label exists, prompts the LLM to propose new Level-1 or Level-2 labels.
+Each insight is represented as a structured 4-tuple: Taxonomy, hierarchical labels for indexing and retrieval; Condition, an explicit description of the applicable prerequisites and triggering signals in the problem; Explanation, the underlying principle of applying this insight; and Example, a concrete demonstration such as corresponding mathematical expressions or code snippet.
+Insight Verification.
+To ensure the retrievability and correctness of newly generated natural-language insights and to prevent the uncontrolled accumulation of idle or invalid entries in the library, each insight undergoes two stages of strict verification before being stored: Retrieval verification: temporarily incorporate the insight into a copy of the library, and check whether the source task can match it via LLM-driven retrieval; Execution verification: reapply the new insight together with other retrieved insights to its source task to verify that the new insight can resolve the original failure and lead to optimality.
+At each stage, a limited number of rewrite attempts is allowed until the insight passes verification, during which the LLM performs targeted revisions of the taxonomy, condition or the whole insight to improve its efficacy.
+Library Storage.
+Experience library is maintained as a dynamically updating taxonomy dictionary organized into three main tracks: Domain Modeling (problem-specific structures and assumptions), General Formulation (reusable mathematical patterns), and Code Implementation (solver-specific coding practices).
+Within each track, the library organizes insights using a two-level labeling scheme, where Level-1 captures a broad category and Level-2 specializes it into a more specific subcategory.
+The taxonomy dictionary is initialized with few-shot labels and expands online: Each new insight is either mapped to an existing category or, if no suitable label exists, prompts the LLM to propose new Level-1 or Level-2 labels.
 Each label is also assigned a condition, written by the LLM, that specifies when the category should be retrieved.
-When storing insights, to reduce redundancy, the LLM also checks whether a similar insight already exists and performs merging when appropriate.
+To prevent uncontrolled growth of the experience library, we incorporate an online merging mechanism throughout training to prevent degradation in retrieval and execution efficiency.
+For each newly generated insight, the system performs redundancy checking with existing insights, merging when appropriate, followed by correctness validation on all associated tasks.
+Library Retrieval.
+During solution generation, retrieved insights from the Domain Modeling and General Formulation tracks guide the construction of the mathematical model, while insights from the Code Implementation track guide solver-code generation.
 To align a target task with relevant insights, we employ a two-step LLM-driven retrieval procedure: Quick label matching, then full applicability check.
 The system first scans the taxonomy dictionary to identify labels that are potentially relevant to the context of the tasks.
-For example, Level-2 label such as Fixed Charge (Big-M Linking) will be probably detected when the problem description specifies that service or flow is allowed only if a facility is opened.
+For example, a Level-2 label such as Fixed Charge (Big-M Linking) will probably be detected when the problem description specifies that service or flow is allowed only if a facility is opened.
 After candidate labels are identified, the system rigorously evaluates each associated insight by examining its condition, and only the most applicable insights are retained.
-During solution generation, retrieved insights from the Domain Modeling and General Formulation tracks guide the construction of the mathematical model, while insights from the Code Implementation track guide solver-code generation.
-This two-step procedure ensures that insights are applied appropriately, while the hierarchical taxonomy provides an extensible structure for organizing and retrieving insights as the library grows.
-Operational Flow.
-Training proceeds in an online regime over minibatches of data, starting from an empty library.
-For each batch, the system retrieves candidate insights, generates and executes programs, and upon failures extracts insights and immediately commits those that pass local self-verification, allowing newly added insights to benefit subsequent tasks and preventing the generation of repetitive insights.
-To reduce redundancy, tasks are clustered and reordered by problem type and semantic similarity, and overlapping insights within a batch are merged prior to integration.
-The process iterates until overall accuracy plateaus, at which point the library is archived and used for evaluation.
-In implementation, for the sake of efficiency, training follows two coordinated data flows.
-The first processes minibatches of tasks in parallel for insight extraction.
-The second maintains a centralized queue of all generated insights, storing them into the library sequentially.
-This queue does not allow asynchronous updates, as concurrent modifications could lead to conflicts if two insights attempt to update the library simultaneously.
-This design balances parallelism in problem-solving with strict serialization in library updates, ensuring both efficiency and consistency.
-3.1.2 Library Evolution
-Figure 4: A locally verified initial condition (light-grey dashed circle) is refined into a broader applicability boundary (solid black) through evolutionary prompt optimization guided by the aggregate metric.
+By employing a progressive disclosure strategy that reveals only task-relevant information when it becomes applicable, this procedure mitigates context overflow while maintaining the effectiveness of retrieval.
+3.2. Library Evolution
 While Library Learning expands the repository of insights, Library Evolution aims to transform task-specific lessons into broadly applicable knowledge.
-Since each insight’s applicability is defined by a condition induced from a specific task, early conditions are often too narrow (failing to trigger on relevant tasks) or too broad (causing misretrieval).
-Left unchecked, these misalignments lead to missed opportunities or systematic failures.
-Library Evolution counters this with a diagnostic–refinement cycle: it detects misaligned insights, aggregates evidence across tasks, and refines conditions at the end of each iteration.
-The refinement is guided by an aggregate metric rather than ad-hox fixes.
+Since each insight’s applicability is defined by a condition induced by a specific task, early conditions are often too narrow (failing to trigger on relevant tasks) or too broad (causing misretrieval).
+Left unchecked, these misalignments lead to widespread misguidance or missed opportunities.
+Library Evolution counters this with a diagnosis–refinement cycle: It detects misaligned insights, aggregates evidence across tasks, and refines conditions at the end of each iteration.
+The refinement is guided by an aggregate metric rather than ad hoc fixes.
 As illustrated in Figure 3, library refinement can be understood as adjusting each insight’s condition toward the correct retrieval boundary in the problem space.
-Diagnostic: Library Diagnosis.
-After each training round, we trace failed tasks and analyze their interaction with the library.
-The diagnostic agent partitions the relationship between each insight $i$ and its associated tasks into three disjoint categories: $\Pi(i)=\{\text{Positive}:S_{i}^{+},\text{Negative}:S_{i}^{-},\text{Unretrieved}:S_{i}^{u}\}$ where $S_{i}^{+}$ contains tasks where the insight was applicable and contributed to the correct formulation, $S_{i}^{-}$ contains tasks where it was misleading and degraded performance, and $S_{i}^{u}$ contains tasks where it was not retrieved but would have been beneficial.
+Library Diagnosis.
+After an initial pass over the training tasks, the system continuously diagnoses task failures and traces their interactions with library insights through iterative cycles of insight retrieval and solution generation across tasks.
+By comparing the formulation and program generated under the guidance of retrieved insights with the ground truth to locate discrepancies, the diagnostic LLM agent partitions the relationship between each insight $i$ and its associated tasks into three disjoint categories: $\Pi(i)=\{\text{Positive}:S_{i}^{+},\text{Negative}:S_{i}^{-},\text{Unretrieved}:S_{i}^{u}\}$ where $S_{i}^{+}$ contains tasks for which the insight was applicable and contributed to the correct formulation and program, $S_{i}^{-}$ contains tasks for which it was misleading and degraded performance, and $S_{i}^{u}$ contains tasks for which it was not retrieved but would have been beneficial.
 By maintaining these partitions across iterations, the system continuously builds a performance profile for each insight.
-If a failed task is subsequently solved after removing a misleading (negative) insight or by injecting a previously unretrieved one, the system attributes the failure to condition misalignment rather than lack of knowledge, thus avoiding redundant insight generation.
-Unretrieved tasks are identified by first comparing the model’s generated program with the ground-truth (or a reference program from self-exploration) to locate discrepancies.
-These discrepancies guide the search for candidate insights, which are then verified for their ability to resolve the errors.
-Verified insights are flagged as relevant but unretrieved, allowing the system to diagnose retrieval gaps without resorting to intractable combinatorial search.
-The left panel of Figure 3 (see the Library diagnosis flow label) illustrates the workflow for this stage.
-Evolver: Library Refinement.
-Building on the diagnosis, the Evolver agent refines insights in two steps: condition refinement and refinement verification.
-First, for each diagnosed insight, the agent strengthens or prunes its applicability condition.
+If a failed task is subsequently solved after removing a misleading (negative) insight or by injecting a previously unretrieved one, the system attributes the failure to condition misalignment rather than lack of knowledge and thus avoids redundant insight generation.
+Figure 2 (Library Diagnosis Flow) illustrates the workflow for this stage.
+Figure 3. A locally verified initial condition is refined into a broader applicability boundary through LLM-driven refinement.
+The figure provides an abstract illustration of how an insight’s applicability condition is refined over time.
+Each point represents a task outcome, where green points indicate tasks for which the insight is helpful, red points indicate tasks for which it is harmful, and blue points indicate tasks where the insight was not retrieved.
+The dashed curve shows the initial applicability condition, which is locally verified but narrow.
+Through library refinement, the condition boundary is adjusted, shown as the solid curve, to better separate positive cases from negative ones and to cover relevant unretrieved cases.
+The arrow indicates the refinement process, resulting in a broader and more accurate applicability boundary for the insight.
+Library Refinement.
+Building on the diagnosed interaction of each insight, the LLM agent first strengthens or prunes its applicability condition.
 Negative tasks contribute explicit inapplicability clauses (e.g., constraints or contexts that block use), while unretrieved tasks highlight missing applicability signals.
-The Evolver then proposes multiple refinement strategies (e.g., adding preconditions, introducing keyword anchors, merging overlapping triggers) and produces candidate conditions with the goal of preserving correct cases, eliminating mismatches, and recovering previously missed tasks.
+Multiple refinement strategies are asked to be proposed (e.g., adding preconditions, introducing keyword anchors, merging overlapping triggers) and produce candidate conditions with the goal of preserving correct cases, eliminating mismatches, and recovering previously missed tasks.
 Then, each candidate condition replaces the original and is tested over the union $R_{i}=S_{i}^{+}\cup S_{i}^{-}\cup S_{i}^{u}$.
@@
 quantifies improvement.
-Here, “kept positives” are tasks that still remain correctly retrieved after refinement; “corrected negatives” are tasks that were misled by the insight before and no longer retrieved; and “recovered unretrieved” are tasks that become correctly retrieved after refinement.
+Here, kept positives are tasks that remain correctly retrieved after refinement; corrected negatives are tasks that were misled by the insight before and are no longer retrieved; and recovered unretrieved are tasks that become correctly retrieved after refinement.
 We accept refinements that increase $p_{i}$ and keep the one with the highest $p_{i}$.
 The workflow for this stage is illustrated in the right panel of Figure 2.
-3.2 Optimization Perspective
+Table 1.
+Accuracy (%) on Test Split and Out-of-Distribution datasets, along with micro- and macro-averaged accuracy across datasets.
+Best per column in bold and second best underlined (ablation variants excluded from highlighting).
+Test Split Out-of-Distribution
+Method NLP4LP (73) NL4OPT (64) IndustryOR (25) MAMO (ComplexLP) (34) Micro Macro LogiOR (92) OptiBench (403) Micro Macro
+Prompt-based Standard 68.5 54.7 52.0 44.1 57.7 54.8 46.7 72.7 67.9 59.7
+OptiMus 71.2 73.4 36.0 29.4 60.2 52.5 17.4 74.7 64.1 46.0
+ORThought 69.9 75.0 60.0 41.2 65.3 61.5 44.6 84.4 77.0 64.5
+Fine-tuning ORLM 86.3 87.5 36.0 55.9 75.0 66.4 19.6 78.2 67.3 48.9
+LLMOPT 97.3 86.5 44.0 85.8 85.0 78.4 40.2 66.4 61.5 53.3
+Experience-learning Reflexion 76.7 64.1 56.0 47.1 64.8 61.0 43.5 76.9 70.7 60.2
+Expel 79.5 67.2 48.0 70.4 69.9 66.3 47.8 80.4 74.3 64.1
+[…]
+Ablation (AlphaOPT) w/o Refinement 82.2 ($\downarrow 4.1$) 80.7 ($\downarrow 0.5$) 44.0 ($\downarrow 16.0$) 61.8 ($\downarrow 14.7$) 73.3 ($\downarrow 6.8$) 67.2 ($\downarrow 9.2$) 48.9 ($\downarrow 7.1$) 92.1 ($\downarrow 1.4$) 84.1 ($\downarrow 2.4$) 70.5 ($\downarrow 4.3$)
+w/o Insight Example 86.3 ($\downarrow 1.4$) 78.1 ($\downarrow 3.1$) 56.0 ($\downarrow 4.0$) 58.8 ($\downarrow 17.7$) 76.0 ($\downarrow 4.1$) 69.8 ($\downarrow 6.6$) 45.7 ($\downarrow 10.3$) 91.0 ($\downarrow 2.5$) 82.9 ($\downarrow 3.6$) 68.4 ($\downarrow 6.4$)
+3.3. Optimization Perspective
 The framework can be viewed as an iterative solution to the optimization problem in the library space.
@@
 The objective is to maximize task success while penalizing library complexity to mitigate retrieval inefficiency and long-context degradation in LLM inference:
-$\max_{\ell\in\mathcal{L}}\;\;\mathbb{E}_{t\sim\mathcal{T}}\!\left[\mathrm{Success}(t\mid\ell)\right]\;-\;\lambda\,\Omega(\ell).$
+$\max_{\ell\in\mathcal{L}}\;\;\mathbb{E}_{t\sim\mathcal{T}}\!\left[\mathrm{Success}(t\mid\ell)\right]\;-\;\lambda\,\Omega(\ell),$
 where $\mathrm{Success}(t\mid\ell)$ indicates whether $\ell$ enables the system to produce a program that achieves the correct optimal objective for task $t$, and $\Omega(\ell)$ quantifies library complexity (e.g., number of insights or redundancy-adjusted size).
 Under our problem design—bounded and continuous property of $\text{Success}(\cdot)$ and $\Omega(\cdot)$, sufficient exploration under solver verification, and bounded merging—the refinement dynamics converge to a locally optimal library.
-In Appendix C, we provide a conceptual sketch showing that convergence holds: As refinement in the second phase strictly improves the first term, while verified merging in the first phase reduces the second term without diminishing the first, sufficient exploration combined with iterative cycles of library learning and evolution ensures convergence to a local optimum.
-Given the inherent ambiguity of natural language and stochasticity in LLM outputs, we present this perspective not as a strict theorem but as a principled justification for the acquisition–refinement design and the redundancy-reduction operations.
-4 EXPERIMENTS
-Our experiments are designed to reflect the requirements that arise in real-world optimization and operations research (OR) applications.
-In these settings, methods are expected not only to perform well on standard benchmarks, but also to transfer across domains, to remain effective when limited supervision is available, to improve steadily as more data becomes available, and to offer results that can be inspected and audited.
-We therefore organize our evaluation around four questions: (1) How well does the method generalize across domains?
-(2) Can it learn effectively with limited supervision?
-(3) Does performance improve consistently with more training data?
-(4) How does it compare overall with strong baselines?
-Finally, we examine the interpretability of the insight library to assess whether the outputs are understandable and actionable to practitioners.
-4.1 Experimental Setup
-Our experiments are conducted on a dataset of 454 problem instances, aggregated from four real-world optimization and operation task datasets, namely the NLP4LP (AhmadiTeshnizi et al., 2024), NL4OPT (Ramamonjison et al., 2022), IndustryOR (Huang et al., 2025), MAMO (ComplexLP) (Huang et al., 2024), with any invalid entries discarded.
-These collections span various formulation types and originate from diverse sources, including academic papers, textbooks, and real‐world industry scenarios.
-Detailed descriptions of these datasets are provided in Appendix B.
+In Appendix A, we provide a conceptual sketch that shows that convergence holds: Since refinement in the second phase strictly improves the first term, while verified merging in the first phase reduces the second term without diminishing the first, sufficient exploration combined with iterative cycles of library learning and evolution ensures convergence to a local optimum.
+Given the inherent ambiguity of natural language and stochasticity in LLM outputs, we present this perspective not as a strict theorem but as a principled justification for the acquisition–refinement design and redundancy-reduction operations.
+4. EXPERIMENTS
+Our experiments are designed to reflect the requirements that arise in real-world optimization and operations research applications.
+In these settings, methods are expected not only to perform well on standard benchmarks but also to transfer across domains, remain effective when limited supervision is available, improve steadily as more data become available, and offer results that can be inspected and audited.
+We organize our empirical evaluation around four questions: (1) How well does the method generalize across domains?
+(2) How effective are the insight content structure and library refinement mechanisms?
+(3) Does performance improve consistently as more training data become available?
+(4) Does the library size converge as training progresses?
+4.1. Experimental Setup
+Our experiments are conducted on a dataset of 452 problem instances, aggregated from four real-world optimization task datasets—the NLP4LP (AhmadiTeshnizi et al., 2024), NL4OPT (Ramamonjison et al., 2022), IndustryOR (Huang et al., 2025), MAMO (ComplexLP) (Huang et al., 2024).
+These collections span various formulation types and originate from diverse sources, which include academic papers, textbooks, and real‐world industry scenarios.
+Detailed descriptions of these datasets are provided in Appendix B.1.
 We perform stratified sampling within each dataset, randomly partitioning $70\%$ for training and $30\%$ for testing.
-We maintain a strict separation between training and test data.
-The experience library is constructed only from training tasks, ensuring that training-derived insight examples do not leak into the test set.
-To assess out-of-distribution generalization, we additionally evaluate on LogiOR (Yang et al., 2025a) and Optibench (Yang et al., 2024).
-Unless otherwise specified, GPT-4o (OpenAI 2024) with temperature 0 is used as the backbone.
+The experience library is constructed only from training tasks, which ensures that training-derived insight examples do not leak into the test set.
+To assess out-of-distribution generalization, we also evaluate on LogiOR (Yang et al., 2025a) and Optibench (Yang et al., 2024).
+Unless otherwise specified, gpt-4o (OpenAI, 2024) with temperature 0 is used as the backbone.
 We use the success rate as the primary evaluation metric, following the evaluation protocol of Yang et al.
-(2025a), where a task is considered successful if the LLM-generated optimal value closely aligns with the provided ground-truth solution.
+(2025a), whereby a task is considered successful if the LLM-generated optimal value closely aligns with the provided ground-truth solution.
+We also report micro- and macro-averaged accuracy for both test-split and out-of-distribution datasets (see Appendix B.2 for definitions).
 Baselines.
-We evaluate against two families of baselines.
-(i) Prompt-based: a vanilla baseline that directly generates the mathematical model from a simple prompt, as well as Reflexion (Shinn et al., 2023), OptiMUS (AhmadiTeshnizi, Gao, and Udell, 2024), and ORThought (Yang et al., 2025a).
-(ii) Learning-based: ORLM (Huang et al., 2025), built on LLaMa3-8B, and LLMOPT (Jiang et al., 2025), built on Qwen2.5-14B (the latest open-source version available after their paper).
-Figure 5: Performance on out-of-distribution datasets. Numbers report test accuracy on LogiOR, OptiBench, and OptMath.
-4.2 Out-of-Distribution Generalization
-We evaluate how well different methods generalize beyond their training distribution.
-For this purpose, we use two benchmarks that were not included during training: LogiOR (Yang et al., 2025a) and OptiBench (Yang et al., 2024).
-Details are provided in Appendix B.
-These datasets were either released after the baseline model (ORLM) or explicitly designed in baseline model’s experiment setting to avoid overlap with their training set (LLMOPT).
-Figure 5 summarizes the results.
-Fine-tuned models such as ORLM and LLMOPT show strong in-distribution performance but exhibit a noticeable drop on unseen datasets.
-For example, ORLM falls to $19.6\%$ on LogiOR and $13.3\%$ on OptMath, while LLMOPT performs better but still degrades compared to its in-distribution performance.
-By contrast, AlphaOPT maintains higher accuracy across all three benchmarks, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench.
-These results support our hypothesis: fine-tuned models tend to learn the syntax of solutions and may perform well when problems are very similar, but they struggle to capture the underlying principles needed for broader problem solving.
-In contrast, the learned experience library equips AlphaOPT with stronger out-of-distribution generalization capability.
-4.3 Learning with Limited Supervision
-In practical OR applications, gold-standard programs are rarely available.
-We therefore test whether AlphaOPT can learn solely from answers.
-Since two datasets in our training set contain gold-standard programs, we remove them in this experiment and let AlphaOPT learn exclusively from answer labels through self-exploration, as introduced in Section 3.1.1.
-As shown in the last two rows of Table 2, remarkably, when trained with answer-only supervision, AlphaOPT achieves accuracy comparable to when it is trained with gold-standard programs.
-AlphaOPT (self-exploration) outperforms all prompt-based methods on test splits of the training data and even achieves the best performance on the OOD OptiBench dataset (92.1% accuracy).
-This demonstrates another advantage of AlphaOPT over fine-tuning–based methods, which require detailed annotations of mathematical formulations and code in order to achieve strong performance.
-Table 1: AlphaOPT steadily improves in both Micro and Macro averages with increasing training size, while maintaining a compact library.
-Training Size MicroAvg MacroAvg Library Size
-100 83.24% 65.80% 38
-200 85.09% 69.22% 103
-300 85.21% 72.12% 110
-4.4 Continual Growth with Data
-We test whether AlphaOPT can improve its performance as more data becomes available.
+We evaluate against three families of baselines.
+(i) Prompt Designing: directly generate the mathematical model from manually crafted prompts, which include OptiMUS (AhmadiTeshnizi et al., 2024), and ORThought (Yang et al., 2025a).
+(ii) Fine-tuning: adapt model parameters through fine-tuning, including ORLM (Huang et al., 2025), built on LLaMa3-8B, and LLMOPT (Jiang et al., 2025), built on Qwen2.5-14B (the latest open-source version available after their paper).
+(iii) Experience Learning: Accumulate reusable knowledge from past task-solving attempts to improve decision quality across tasks, including two representative works Reflexion (Shinn et al., 2023) and ExpeL (Zhao et al., 2024).
+For ExpeL, we adapt the original framework to operations research tasks by making minor changes to task-related prompts and evaluation criteria.
+4.2. Overall Performance
+Table 2 summarizes the overall performance across test and out-of-distribution benchmarks.
+AlphaOPT consistently outperforms prompt-based and experience-learning baselines on the test splits and achieves the best performance on the more challenging LogiOR, IndustryOR, and out-of-distribution OptiBench datasets, exceeding the second-best method by an average of 7%.
+Although fine-tuned models (ORLM and LLMOPT) achieve higher scores on certain in-distribution benchmarks such as NLP4LP and NL4OPT, these gains are less conclusive, as their training data either overlap with our evaluation splits or are not publicly available.
+Moreover, many existing benchmarks are derived from a small set of seed problems (Ramamonjison et al., 2022; Huang et al., 2024), which favors fine-tuning approaches that excel at pattern memorization.
+In contrast, while fine-tuning methods require detailed annotations of mathematical formulations and code to achieve strong performance, AlphaOPT attains competitive results using only answer labels.
+This advantage is particularly significant in OR practice, where high-quality, expert-annotated datasets are generally scarce.
+Compared with baseline experience-learning methods, our approach achieves superior performance; our analysis in Appendix B.3 further shows that methods such as ExpeL tend to learn generic insights that lack the detailed modeling and coding guidance needed to prevent common errors.
+In particular, we evaluate robustness to distribution shift using two benchmarks not seen during training: LogiOR (Yang et al., 2025a) and OptiBench (Yang et al., 2024).
+Fine-tuned models show strong in-distribution performance but degrade substantially on these unseen datasets.
+For example, ORLM drops to $19.6\%$ accuracy on LogiOR and $13.3\%$ on OptiBench, while LLMOPT also exhibits noticeable degradation.
+In contrast, AlphaOPT maintains substantially higher accuracy, achieving $56.0\%$ on LogiOR and $93.5\%$ on OptiBench.
+These results suggest that fine-tuning tends to capture dataset-specific solution patterns, whereas the experience library in AlphaOPT enables the learning of transferable modeling principles, leading to stronger robustness under distribution shift.
+4.3. Ablation Study
+We examine two ablations of AlphaOPT: removing the library refinement stage (w/o Refinement) and removing concrete examples from insights (w/o Insight Example).
+In w/o Refinement, insights are stored after extraction without further adjustment of their applicability conditions.
+In w/o Insight Example, insights exclude concrete mathematical or code-level demonstrations.
+Removing refinement leads to consistent performance drops across both test-split and out-of-distribution datasets.
+On the test split, micro- and macro-averaged accuracy decrease by 6.8 and 9.2 points respectively, with the largest degradation on structurally complex datasets such as IndustryOR and MAMO (ComplexLP).
+Out-of-distribution performance also declines, indicating that unrefined applicability conditions are insufficient for reliable cross-task reuse.
+Removing insight examples results in a smaller but systematic drop in aggregate accuracy, with a notably larger impact on formulation-to-code–heavy datasets.
+4.4. Continual Growth and Library Converge
+We test whether AlphaOPT can improve its performance as more data become available.
 We incrementally sample sets of 100, 200, and 300 data items from our training set and train AlphaOPT on each subset.
 As shown in Table 2, when evaluated on out-of-distribution datasets (LogiOR, OptiBench), we observe that AlphaOPT steadily improves its performance with increasing data size, without requiring updates to its model parameters.
-4.5 Overall Performance
-AlphaOPT achieves the best accuracy on out-of-distribution datasets, reaching $51.1\%$ on LogiOR and $91.8\%$ on OptiBench.
-On in-distribution test splits, fine-tuned models such as ORLM and LLMOPT achieve higher scores on certain datasets (e.g., LLMOPT obtains $97.3\%$ on NLP4LP and $85.8\%$ on MAMO).
-However, these advantages are less conclusive, since LLMOPT’s training data are not publicly available and may overlap with our test splits.
-Moreover, many existing benchmarks are derived from a small set of seed problems (Ramamonjison et al., 2022; Huang et al., 2024), which favors fine-tuning approaches that excel at pattern memorization.
-In contrast, AlphaOPT performs competitively across all in-distribution datasets, matches or exceeds baselines on IndustryOR and MAMO(ComplexLP), and maintains a clear margin on out-of-distribution generalization.
-These results demonstrate that the experience library enables AlphaOPT to learn transferable modeling principles rather than dataset-specific syntax, resulting in stronger robustness to distribution shifts.
-Table 2: Accuracy on in-distribution Test Split and Out-of-Distribution datasets (higher is better). Best per column in bold.
-Test Split Out-of-Distribution
-Method
-NLP4LP
-NL4OPT
-IndustryOR
-MAMO
-(ComplexLP) (34)
-LogiOR
-OptiBench
-Prompt-based Standard $68.5$ $54.7$ $52.0$ $44.1$ $46.7$ $72.7$
-Reflexion $76.7$ $64.1$ $56.0$ $47.1$ $43.5$ $76.9$
-OptiMus $71.2$ $73.4$ $36.0$ $29.4$ $17.4$ $74.7$
-ORThought $69.9$ $75.0$ 60.0 $41.2$ $44.6$ $84.4$
-Fine-tuning-based ORLM $86.3$ 87.5 $36.0$ $55.9$ $19.6$ $78.2$
-LLMOPT 97.3 $86.5$ $44.0$ 85.8 $40.2$ $66.4$
-Ours
-AlphaOPT
-$83.6$ $79.7$ 60.0 $85.3$ 51.1 $91.8$
-AlphaOPT
-(self-exploration)
-$86.3$ $79.7$ 60.0 $76.5$ $50.0$ 92.1
-5 Library Analysis
-To fully interpret the experience library, we visualize its structure and the distribution of insights under a hierarchical taxonomy.
-Additionally, we select several representative insights to analyze the effectiveness of library refinement, which are provided in the Appendix D.3.
-Figure 6: The learned experience-library taxonomy has three hierarchy levels: 3 main tracks, 14 level-1 labels, and 38 level-2 labels.
-This figure presents the level-1 and level-2 terms and their assignments.
-Percentages in red, shown in parentheses after each track and level-1 term, denote that category’s percentage within its parent.
-The detailed specification and distribution of level-2 labels are provided in the Appendix D.2 and Appendix D.4.
-As illustrated in Figure 6,
-1) The library mainly captures insights that address LLMs’ difficulties with domain-specific modeling structures, particularly those involving structural coupling and constraint balance.
-In the library taxonomy, the number of insights under Domain Modeling track accounts for 52% of the total.
-Within this track, the most frequent level-1 labels are Resource Allocation (38.9%), Network Flow (31.9%), and Production Planning (12.5%).
-Structural coupling, which refers to the model’s difficulty in capturing cross-variable or cross-stage dependencies, is reflected in level-2 labels such as Fixed Charge (Big-M Linking), Set-Partitioning/Assignment, and Time-Indexed Variables.
-Constraint balance, which refers to the model’s failure to maintain system-wide conservation and resource equilibrium, is reflected in level-2 labels such as Capacity/Resource Balance Equations, Flow Conservation, and Inventory Balance Equations.
-2) The library captures insights that help transform intuitive or context-based reasoning into mathematically rigorous and solver-consistent formulations, particularly in defining variables, formalizing constraints, and maintaining numerical consistency.
-General Formulation track accounts for 30% of all insights.
-Within this track, the the most frequent level-1 labels are Variable Definition (66.7%), followed by Constraint Formulation (16.7%) and Units Scaling (11.9%).
-Variable definition difficulty lies in specifying variable domains and maintaining structural consistency, as reflected in level-2 labels such as Explicit Bounds, Continuous vs.
-Discrete Confusion, and Indexing & Data Structure Alignment.
-Constraint formulation difficulty lies in the model’s inaccuracy in representing logical and numerical relationships, as reflected in level-2 labels such as Incorrect Relational Operators, Big-M & Indicator Variables, and Linearization of Non-linear Forms.
-Units scaling difficulty lies in maintaining numerical coherence and stability, as reflected in level-2 labels such as Unit Inconsistency, Big-M Magnitude & Numerical Stability.
-3) The library captures insights that help bridge the gap between symbolic formulations and executable solver code, particularly in handling solver syntax and maintaining data consistency.
-Code Implementation track accounts for 18% of all insights, this proportion is lower than formulation-level insights though, it still exposes critical weaknesses in the execution stage.
-Within this track, most errors originate from Solver & API Syntax (76.0%), while the remaining issues arise from Data I/O & Validation (24.0%).
-Solver & API syntax errors reflect the model’s lack of solver awareness and insufficient control over formal expression generation, as evidenced by level-2 labels such as Nonlinear Constraint Handling, Strict Inequalities, Quicksum vs.
-sum Syntax.
-Data I/O & validation errors, on the other hand, reveal instability in aligning mathematical indices with data structures, as reflected in level-2 labels such as KeyError & Index Mismatch and Matrix Dimension Mismatch.
-6 Success And Failure Case Study
-To demonstrate the effectiveness of the retrieved library insights, we perform case studies across all evaluation datasets (including the test splits of NLP4OPT, NLP4LP, IndustryOR, and MAMO-complex, as well as the out-of-distribution sets LogiOR and Optibench), comparing the performance of solving optimization tasks with and without library retrieval enabled, while excluding tasks that succeed in both settings from the evaluation.
-As shown in Figure 7, approximately half of the retrieved insights successfully contributed to solving new optimization tasks, with only a small portion introducing new errors (Failure), demonstrating the effectiveness of our method in automating knowledge transfer for optimization modeling.
-Moreover, insights under different level-1 taxonomy labels exhibit distinct performance patterns across modeling dimensions.
-(a) Overall Effectiveness of Retrieved Library Insights
-(b) Effectiveness Distribution across Level-1 Insight Categories
-Figure 7: We evaluate the effectiveness of retrieved insights by classifying their outcomes as success, failure, or invalid.
-An insight is considered successful if it matches a new task and prevents an error that would otherwise occur, failed if it mismatches a task and introduces a new error, and invalid if it matches correctly but fails to help the LLM fix the original mistake.
-(a) The overall distribution of insight outcomes.
-(b) The proportions of successful, failed, and invalid insights under level-1 taxonomy labels.
-1) Insights with high success rates
-These insights are concentrated in Code Implementation – Solver & API Syntax and General Formulation – Variable Definition.
-These insights typically target code implementation or fundamental modeling errors with clear structures, enabling the LLM to follow their guidance stably and correctly.
-For instance, insights tagged with Strict Inequality label under Solver & API Syntax highlights that solvers (e.g., Gurobi) do not support strict inequalities and should instead be reformulated as non-strict forms ($\leq-1$ or $\geq+1$); insights tagged with Explicit Bounds label under Variable Definition emphasizes that decision variables should be assigned explicit upper and lower bounds to ensure feasibility and improve solver efficiency.
-2) Insights with high failure rates
-These insights are mainly found in Domain Modeling – Facility Location.
-They often involve structural constraints and logical triggers that are easily misinterpreted or overgeneralized by the LLM.
-For example, insights tagged with Fixed Charge (Big-M Linking) label exhibit a failure rate of 70%.
-Although the principle of using Big-M constraints to model on/off logic is correct, its blind application in problems without conditional activation can lead to redundant or overlapping constraints and unnecessary feasible-region reduction, ultimately causing solver failure.
-Additionally, while insights under Variable Definition generally achieves high success rates, some insights tagged with Explicit Bounds label sometimes are overly rigid, leading the LLM to impose unnecessary upper bounds, thereby restricting the feasible space and producing suboptimal solutions.
-3) Insights with high invalid rates
-These insights are mainly under Domain Modeling – Resource Allocation and Domain Modeling – Product Planning.
-Although the LLM successfully retrieves the correct insights and identifies the corresponding problem types, it often fails to translate them into executable formulas or solver-level implementations.
-For instance, under Solver & API Syntax, Nonlinear Constraint Handling advises linearization or the introduction of auxiliary variables for nonlinear objectives or constraints (e.g., ratios or divisions), yet the LLM frequently fails to fully execute these transformations (neglecting auxiliary variables or mis-rewriting proportional constraints), resulting in insights being recognized but not operationalized.
-Additionally, certain tasks remained unsolved regardless of whether library retrieval was enabled, as their failure stemmed from factors beyond the current learned library’s knowledge scope.
-In the out-of-distribution dataset LogiOR, tasks involved multi-level spatiotemporal logic and interacting constraints (e.g., capacity, timing, and flow balance) in problems such as routing, scheduling, and network flow.
-These challenges extend beyond the scope of the existing library, which primarily focuses on static, linear formulations.
-Although related taxonomy labels such as Resource Allocation and Nonlinear Constraint Handling exist, their granularity and depth remain insufficient for modeling such complex logic, revealing that the current system, while semantically generalizable, still lacks robust cross-structural transfer and context adaptation capabilities.
-Overall, while successful insights constitute the majority, the results reveal several directions for improvement.
-First, the relatively high failure rates indicate that, despite the inclusion of condition refinement, retrieval precision can still be improved through enhanced semantic disambiguation and structural filtering.
-Second, insights with high invalid proportions suggest the need for clearer explanations and better-designed examples to improve pedagogical clarity and execution effectiveness.
-Finally, for out-of-distribution tasks, future efforts should focus on strengthening the LLM’s ability to adapt and generalize retrieved insights to unseen, complex optimization scenarios.
-Moreover, expanding OR datasets based on LLM error typologies can further enhance experiential learning efficiency and generalization at comparable problem scales.
-7 Conclusion and Discussion
+Also, we provide empirical validation of the theoretical analysis presented in Section 3.3, under which AlphaOPT is expected to converge to a locally optimal experience library during training.
+As illustrated in Figure 4, the cumulative training accuracy increases rapidly with the number of visited tasks in the early stage and gradually saturates thereafter.
+Meanwhile, library size gradually slows and eventually stabilizes as training proceeds.
+This behavior results from the combined effects of online insight merging, solver-based verification, and the refinement mechanism, under which the number of distinct and verifiable insights that can be incorporated in the library progressively saturates.
+Taken together, the joint observation of continual performance improvement and library convergence indicates that AlphaOPT achieves scalable learning without unbounded accumulation of redundant or marginally useful insights.
+Table 2. AlphaOPT steadily improves in both Micro and Macro averages with increasing training size.
+Training Size MicroAvg (%) MacroAvg (%)
+100 83.24 65.80
+200 85.09 69.22
+300 85.21 72.12
+Figure 4. Training dynamics of the experience library. As training progresses, accuracy converges while library growth stabilizes.
+The figure shows the training dynamics of the experience library across iterations.
+The left plot reports cumulative accuracy as more tasks are visited.
+Accuracy increases rapidly at the beginning, then improves more gradually, and eventually stabilizes as additional tasks are processed.
+The right plot shows the size of the experience library over training iterations.
+The library grows steadily in the early iterations and then remains mostly the same after four iterations.
+Together, the two plots illustrate that performance converges while the library size stabilizes.
+5. Library Analysis
+5.1. Library Taxonomy and Knowledge
+Figure 5. The learned experience library taxonomy consists of three main tracks, 14 level-1 labels, and 38 level-2 labels. Percentages shown in red indicate each track’s or level-1 label’s proportion.
+The figure shows the structure of the learned experience library taxonomy.
+The taxonomy is organized into three main tracks: Domain Modeling, General Formulation, and Code Implementation.
+Each track is further divided into level-1 categories, which are then broken down into more specific level-2 concepts.
+Domain Modeling covers common problem structures such as resource allocation, network flow, production planning, balancing, packaging, facility location, selection, and routing.
+For each category, representative modeling elements are listed, such as capacity and flow balance equations, fixed-charge modeling, inventory balance, and subtour elimination.
+General Formulation focuses on how optimization models are expressed, including variable definition, constraint formulation, unit scaling, and objective specification, with examples of typical pitfalls such as continuous versus discrete confusion, Big-M misuse, and incorrect objective definitions.
+Code Implementation captures issues that arise when translating formulations into solver code, including solver and API syntax, constraint expression syntax, and data input, output, and validation errors.
+Percentages shown in red indicate the proportion of experiences assigned to each track and to each level-1 category, reflecting how frequently different types of insights appear in the learned library.
+To fully interpret the experience library, we visualize its hierarchical taxonomy and the distribution of insights (as shown in Figure 5).
+Detailed specification and distribution of level-2 labels are provided in the Appendix B.5 and B.7.
+Overall, LLMs tend to make mistakes more frequently on mathematical modeling than in code implementation when solving optimization problems, especially when problem structures are complex or the modeling assumptions and conventions are implicit in the context.
+Specifically, challenges in LLM-driven optimization can be summarized as threefold:
+Structural Coupling and Balance.
+In the Domain Modeling track, LLMs exhibit two primary difficulties: (i) capturing structural coupling across variables or stages, and (ii) maintaining global balance constraints over system resources or quantities.
+These challenges are particularly pronounced in problem classes such as Resource Allocation (38.9%), Network Flow (31.9%), and Production Planning (12.5%).
+For example, in production planning scenarios, LLMs often fail to correctly couple machine-level production decisions with inter-temporal inventory balance constraints; they rely instead on aggregated variables to constrain total processing time, and thus violate essential structural dependencies.
+From Intuition to Rigorous Formulation.
+In the General Formulation track, LLMs are prone to pitfalls when translating intuitive or contextual descriptions into mathematically rigorous and solver-consistent formulations, with mistakes concentrated in Variable Definition (66.7%), Constraint Formulation (16.7%), and Unit Scaling (11.9%).
+These errors include unclear variable domains, confusion between continuous and discrete variables, misuse of relational operators, improper handling of Big-M or indicator variables, and unit inconsistencies.
+For example, in production or resource constraints expressed as at least or at most (e.g., minimum production requirements or budget limits), LLMs often model such conditions as equalities (==) or strict inequalities (<< or >>), leading to overly restrictive formulations.
+Bridging Models and Executable Code.
+In the Code Implementation track, although such errors are less frequent, LLLMs still exhibit gaps when translating symbolic mathematical models into executable solver code.
+Among these errors, 76% stem from Solver & API Syntax (e.g., improper handling of nonlinear constraints, confusion between quicksum and sum), while the remaining 24% arise from Data I/O & Validation issues (e.g., KeyError, index mismatches, and matrix dimension errors).
+5.2. Examples of Library Refinement
+Representative refined insights are selected from the library to illustrate how their applicability conditions are adjusted based on the associated tasks.
+Figure 6. Examples of insight applicability conditions before and after refinement, along with the corresponding changes in task associations.
+The figure shows concrete examples of insight applicability conditions before and after refinement, together with how task associations change as a result.
+The figure presents three representative insights, each identified by an ID, along with their original applicability conditions and the refined conditions learned through library refinement.
+The examples cover different modeling patterns, including single-capacity item selection, resource flow modeling with conservation constraints, and max-flow formulations between a source and a sink.
+For each insight, the refined condition is more specific about the problem features that must be present and explicitly states when the insight should not be applied.
+At the bottom right, the figure shows a bar chart summarizing how tasks are associated with each insight before and after refinement.
+Task outcomes are grouped into positive, negative, and unretrieved cases for both the original and refined conditions.
+After refinement, positive associations increase while negative and unretrieved associations decrease, indicating that the refined conditions better match the tasks where each insight is appropriate.
+As shown in the Figure 6, Insight ID 10 targets single–scalar-capacity selection; the refinement broadens coverage by adding equivalent formulations such as minimizing a single resource (time/space), while explicitly excluding multi-capacity settings.
+Insight ID 43 captures flow/conservation and production balancing across nodes; refinement adds lexical anchors (e.g., capacity constraints, balancing production) and excludes statements dominated by fixed proportions, quality constraints, or predefined allocations, and gating suppresses spurious retrievals whereby the structure is not true flow/conservation.
+Insight ID 133 focuses on max-flow between a single source and a single sink; refinement tightens the structural requirement and rules out source-to-many distribution tasks.
+Across the cases, four LLM-driven refinement strategies recur: (i) generalization via equivalent phrasings; (ii) lexical anchoring with positive keywords; (iii) explicit exclusions to reduce misalignment; and (iv) structural qualifiers to prevent misuse.
+As shown in Figure 6, these semantic refinements reduce both negative and unretrieved cases while preserving most positive cases.
+6. Conclusion and Discussion
 This paper addresses the limitations of previous methods by presenting a novel self-improving library learning framework, AlphaOPT, for formulating optimization programs.
@@
 The learned experience library reveals LLMs’ characteristic failure patterns and underlying causes across domain-specific modeling, mathematical formulation, and solver syntax handling.
-Case study results show that insights with high success rate primarily address fundamental modeling errors with clear structures, while insights with high mismatch rate or low effectiveness expose the overgeneralization of structurally complex knowledge and highlight the need further to improve the pedagogical clarity and applicability conditions of insights.
+Case study results show that insights with a high success rate primarily address fundamental modeling errors with clear structures, while insights with a high mismatch rate or low effectiveness expose the overgeneralization of structurally complex knowledge and highlight the need to improve the pedagogical clarity and applicability conditions of insights.
 Looking ahead, we highlight three promising directions.
-First, reasoning-oriented test-time scaling, which is already powerful in other domains, could be particularly effective for OR formulations, where results are inherently verifiable.
-Second, strengthening datasets with both academic research and large-scale real-world industry problems will move LLM systems beyond the toy examples that dominate current benchmarks, enabling progress toward truly large-scale optimization tasks.
-Third, moving beyond correctness toward improving the efficiency of formulations is crucial for real-world deployment, and our self-improving library learning approach offers a promising path toward that goal.
-Appendix
-Appendix A Comparative Analysis of AlphaOPT against Prior Experience-Learning Methods
-Recent approaches such as Reflexion (Shinn et al., 2023), STaR (Zelikman et al., 2022), ExpeL (Zhao et al., 2024), and AlphaEvolve (Novikov et al., 2025) demonstrate that large models can improve through experiential reuse, storing reflections, rationales, or code edits and applying them in new tasks.
-These methods have been effective in open-ended reasoning and programming, but they face limitations for optimization problems.
-First, their experiences are largely preserved as free-form text or edits without explicit applicability semantics, yet in optimization tasks, applying such experiences inappropriately can have detrimental effects.
-Second, their verification is limited to task outcomes such as checking rewards, final answers, or passing test cases, which does not guarantee that the underlying knowledge is structurally valid or transferable.
-Our framework adapts experience learning to operations research (OR) with three key innovations: (1) solver-guided verifiability: correctness is judged at the program level.
-If a program achieves the optimal objective under the solver, it is highly likely to be valid and can serve as a reliable anchor for extracting insights, broadening the sources of experience collection.
-New and refined insights are explicitly re-tested on associated tasks, ensuring they are valid before integration; (2) structured knowledge for interpretability and auditability: each insight is represented with taxonomy, condition, explanation, and example, making its applicability explicit, reviewable, and even revisable in practice; (3) refinement of experience applicability for generalizability and preciseness: applicability conditions are refined using cross-task evidence, so insights neither over-generalize nor remain too narrow, improving safe transfer across problem families.
-See Table  3 for detailed comparisons.
-Table 3: Comparison of experience-learning methods.
-Prior works improve through experiential reuse but rely on free-form knowledge and outcome-level verification.
-Our framework introduces structured insights, solver-guided verification, and refined applicability, which are crucial for OR.
-[…]
-Reflexion Reward signal General agents
-STaR Answer correctness QA / reasoning
-ExpeL (✓) minimal Task success assumed General agents
-AlphaEvolve (✓) implicit Test harness (partial) Code synthesis / evolution
-AlphaOPT Solver optimality + insight verification OR formulation and solver code
-Appendix B Datasets
-We have collected the publicly available optimization problem datasets listed in the Table  4, which include both natural language problem descriptions and their optimal solutions.
+First, reasoning-oriented test-time scaling, which is already powerful in other domains, could be particularly effective for OR formulations, whose results are inherently verifiable.
+Second, strengthening datasets with both academic research and large-scale real-world industry problems will move LLM systems beyond the toy examples that dominate current benchmarks, and thus enable progress toward truly large-scale optimization tasks.
+Third, moving beyond correctness toward improving the efficiency of formulations is crucial for real-world deployment, and our self-improving library learning approach offers a promising path to that goal.
+Acknowledgements.
+To Robert, for the bagels and explaining CMYK and color spaces.
+Appendix A Proof of the Library Convergence
+Recall the optimization problem in the library training phase
+$F(\ell)\;=\;\mathbb{E}_{t\sim\mathcal{T}_{\mathrm{train}}}\!\big[\,r(t\mid\ell)\,\big]\;-\;\lambda\,\Omega(\ell),$
+where $r(t,\ell)$ is a bounded reward function that implements the role of the original $\mathrm{Success}(t\mid\ell)$ (i.e., it measures the matching quality between optimization problem $t$ and library $\ell$), and $\Omega(\ell)$ is a bounded complexity penalty.
+According to the problem setting, the iterative refinement algorithm satisfies:
+  1. (1)
+    Monotone update: At iteration $k$, from $\ell_{k}$, the algorithm considers a set of admissible refinements $R(\ell_{k})\subseteq\mathcal{L}$.
+Each accepted iteration consists of one of two types of operations:
+    • Merge step: decreases $\Omega(\ell)$ while leaving $r(t,\ell)$ non-decrease for all relevant tasks;
+    • Exploration step: improves $r(t,\ell)$ for some tasks without increasing $\Omega(\ell)$.
+    Therefore, every accepted refinement strictly increases $F(\ell)$; otherwise, the algorithm keeps $\ell_{k+1}=\ell_{k}$.
+  2. (2)
+    Sufficient exploration: Any improving neighbor $\tilde{\ell}\in R(\ell_{k})$ (i.e.
+one with a strictly larger objective) will eventually be discovered and executed.
+Empirically, this is achieved through iterative prompt optimization with LLMs.
+  3. (3)
+    Boundedness: $r(t,\ell)$ and $\Omega(\ell)$ are bounded; hence $F(\ell)$ is bounded above and below.
+The following theorem establishes that, under the assumption that the training and testing distributions are identical, the refinement procedure yields libraries that are locally optimal for the testing objective.
+Theorem A.1.
+Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$.
+If the library space $\mathcal{L}$ is finite, the algorithm terminates in finitely many steps at a library $\ell^{*}$ that is a local maximizer for the testing objective.
+Moreover, the algorithm cannot terminate at a saddle point.
+Proof.
+Every accepted merge or exploration step strictly increases $F(\ell)$; otherwise, the library remains unchanged.
+Since $F$ is bounded above, the sequence $\{F(\ell_{k})\}$ is monotone non-decreasing and bounded, and hence convergent to some limit $F^{*}$.
+Furthermore, since $\mathcal{L}$ is finite, define
+$\delta\;=\;\min\big\{F(\tilde{\ell})-F(\ell):\tilde{\ell}\in R(\ell),\;F(\tilde{\ell})>F(\ell)\big\}.$
+Finiteness guarantees $\delta>0$, so only finitely many strict improvements are possible.
+The algorithm halts at some $\ell^{*}$.
+By sufficient exploration, no improving neighbor of $\ell^{*}$ exists.
+Therefore, $\ell^{*}$ is a local maximizer for both training and testing objectives.
+Saddle points are excluded.
+Since the training and testing distributions coincide, the training objective equals the testing objective; thus any local optimality statement directly applies to testing. ∎
+Although the assumption of a finite library is reasonable, we also provide a proof for the case of an infinite library for completeness and rigor.
+Theorem A.2 (Infinite compact library case).
+Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$.
+If the library space $\mathcal{L}$ is compact (closed and bounded) and $F$ is continuous, then the sequence $\{F(\ell_{k})\}$ converges, and any subsequential limit point $\ell^{\infty}$ is a local maximizer for the testing objective.
+Saddle points are excluded for all such limit points.
+Proof.
+Each accepted step strictly increases $F(\ell)$, so $\{F(\ell_{k})\}$ is monotone non-decreasing.
+Since $F$ is bounded above, $\{F(\ell_{k})\}$ converges to some $F^{*}$.
+By compactness of $\mathcal{L}$, there exists a convergent subsequence $\ell_{k_{j}}\to\ell^{\infty}$.
+The continuity of $F$ ensures $F(\ell_{k_{j}})\to F(\ell^{\infty})=F^{*}$.
+Suppose $\ell^{\infty}$ had a neighbor $\tilde{\ell}\in R(\ell^{\infty})$ with $F(\tilde{\ell})>F(\ell^{\infty})$.
+Then sufficient exploration would eventually yield $F(\ell_{k})>F^{*}$, which is a contradiction.
+Therefore, $\ell^{\infty}$ is a local maximizer.
+Saddle points are excluded.
+Since the training and testing distributions coincide, the training objective equals the testing objective; thus any local optimality statement directly applies to testing.
+∎
+Theorems A.1 and A.2 together guarantee that, when the training and testing distributions coincide, the refinement algorithm converges to locally optimal solutions for the testing phase.
+Appendix B Experimental Details
+B.1. Datasets
+We have collected the publicly available optimization problem datasets listed in Table  3, which include both natural language problem descriptions and their optimal solutions.
 Among them, NLP4LP, NL4OPT, IndustryOR, and MAMO (ComplexLP) are used to construct our training and test datasets, while Optibench and LogiOR serve as out-of-distribution (OOD) evaluation sets.
@@
 Accordingly, we train and evaluate on clean splits that exclude instances labeled as erroneous, and the Size column in Table 3 indicates the size of each dataset after cleaning.
-Specifically, for NLP4LP, IndustryOR and LogiOR we use the cleaned versions provided by Yang et al.
-(2025a); for NL4OPT, MAMO (ComplexLP) and Optibench we use the cleaned releases from Astorga et al.
+Specifically, for NLP4LP, IndustryOR, and LogiOR we use the cleaned versions provided by Yang et al.
+(2025a); for NL4OPT, MAMO (ComplexLP), and Optibench, we use the cleaned releases from Astorga et al.
 (2025), obtained from their GitHub repository.
-Table 4: The statistics of the optimization problem datasets
+Table 3. Statistics of the optimization problem datasets
 Dataset Size Formulation Type(s) Completion
@@
 Abbreviations: LP – Linear Programming; IP - Integer Programming; NLP – Nonlinear Programming; MI – Mixed-Integer; others - Quadratic Programming, Dynamic&Stochastic Programming, etc.
-Appendix C Proof of the Library Convergence
-Recall the optimization problem in the library training phase
-$F(\ell)\;=\;\mathbb{E}_{t\sim\mathcal{T}_{\mathrm{train}}}\!\big[\,r(t\mid\ell)\,\big]\;-\;\lambda\,\Omega(\ell),$
-where $r(t,\ell)$ is a bounded reward function that implements the role of the original $\mathrm{Success}(t\mid\ell)$ (i.e., it measures the matching quality between optimization problem $t$ and library $\ell$), and $\Omega(\ell)$ is a bounded complexity penalty.
-According to the problem setting, the iterative refinement algorithm satisfies:
-  1. 1.
-    Monotone update: At iteration $k$, from $\ell_{k}$, the algorithm considers a set of admissible refinements $R(\ell_{k})\subseteq\mathcal{L}$.
-Each accepted iteration consists of one of two types of operations:
-    • Merge step: decreases $\Omega(\ell)$ while leaving $r(t,\ell)$ non-decrease for all relevant tasks;
-    • Exploration step: improves $r(t,\ell)$ for some tasks without increasing $\Omega(\ell)$.
-    Therefore every accepted refinement strictly increases $F(\ell)$; otherwise the algorithm keeps $\ell_{k+1}=\ell_{k}$.
-  2. 2.
-    Sufficient exploration: Any improving neighbor $\tilde{\ell}\in R(\ell_{k})$ (i.e.
-one with strictly larger objective) will eventually be discovered and executed.
-Empirically, this is achieved through iterative prompt optimization with LLMs.
-  3. 3.
-    Boundedness: $r(t,\ell)$ and $\Omega(\ell)$ are bounded, hence $F(\ell)$ is bounded above and below.
-The following theorem establishes that, under the assumption that the training and testing distributions are identical, the refinement procedure yields libraries that are locally optimal for the testing objective.
-Theorem 1.
-Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$.
-If the library space $\mathcal{L}$ is finite, then the algorithm terminates in finitely many steps at a library $\ell^{*}$ which is a local maximizer for the testing objective.
-Moreover, the algorithm cannot terminate at a saddle point.
-Proof.
-Every accepted merge or exploration step strictly increases $F(\ell)$, and otherwise the library remains unchanged.
-Since $F$ is bounded above, the sequence $\{F(\ell_{k})\}$ is monotone non-decreasing and bounded, hence convergent to some limit $F^{*}$.
-Furthermore, since $\mathcal{L}$ is finite, define
-$\delta\;=\;\min\big\{F(\tilde{\ell})-F(\ell):\tilde{\ell}\in R(\ell),\;F(\tilde{\ell})>F(\ell)\big\}.$
-Finiteness guarantees $\delta>0$, so only finitely many strict improvements are possible.
-The algorithm halts at some $\ell^{*}$.
-By sufficient exploration, no improving neighbor of $\ell^{*}$ exists.
-Therefore, $\ell^{*}$ is a local maximizer for both training and testing objectives.
-Saddle points are excluded.
-Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing. ∎
-Although the assumption of a finite library is reasonable, we also provide a proof for the case of an infinite library for completeness and rigor.
-Theorem 2 (Infinite compact library case).
-Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$.
-If the library space $\mathcal{L}$ is compact (closed and bounded) and $F$ is continuous, then the sequence $\{F(\ell_{k})\}$ converges, and any subsequential limit point $\ell^{\infty}$ is a local maximizer for the testing objective.
-Saddle points are excluded for all such limit points.
-Proof.
-Each accepted step strictly increases $F(\ell)$, so $\{F(\ell_{k})\}$ is monotone non-decreasing.
-Since $F$ is bounded above, $\{F(\ell_{k})\}$ converges to some $F^{*}$.
-By compactness of $\mathcal{L}$, there exists a convergent subsequence $\ell_{k_{j}}\to\ell^{\infty}$.
-Continuity of $F$ ensures $F(\ell_{k_{j}})\to F(\ell^{\infty})=F^{*}$.
-Suppose $\ell^{\infty}$ had a neighbor $\tilde{\ell}\in R(\ell^{\infty})$ with $F(\tilde{\ell})>F(\ell^{\infty})$.
-Then sufficient exploration would eventually yield $F(\ell_{k})>F^{*}$, which is a contradiction.
-Therefore, $\ell^{\infty}$ is a local maximizer.
-Saddle points are excluded.
-Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing.
-∎
-Theorems 1 and 2 together guarantee that, when the training and testing distributions coincide, the refinement algorithm converges to locally optimal solutions for the testing phase.
-Appendix D Additional Results
-D.1 Ablation Study
-We assess the effectiveness of library insight retrieval/application and self-debug:
-  • w/o self-debug: Remove the model’s self-debugging section.
-  • w/o taxonomy: Remove the library taxonomy and directly match insights by checking all conditions.
-  • w/o insight example: Use only the explanation as input, excluding exemplar cases
-Table 5 shows that the full AlphaOPT achieves the best scores on all out-of-distribution datasets.
-Removing self-debug yields the largest drop on Logior (15.21%), indicating it plays an important role in the system.
-Dropping the library taxonomy reduces accuracy by 4.34% on Logior and by 1.09% on Optibench, suggesting that structured matching meaningfully improves retrieval.
-Excluding insight examples similarly lowers performance, showing that concrete, worked snippets aid application beyond textual explanations alone.
-Table 5: The ablation results of AlphaOPT without: performance on benchmarks.
-Dataset
-Logior (92)
-Optibench (403)
-AlphaOPT (full) 51.08% 91.81%
-w/o self-debug 35.87% 89.26%
-w/o taxonomy 46.74% 90.72%
-w/o insight example 45.65% 91.06%
-D.2 Details of Library Insight Distribution
+B.2. Micro- and Macro-Averaged Accuracy
+Let $\mathcal{D}=\{D_{1},\dots,D_{K}\}$ be the set of evaluation datasets, with $|D_{k}|$ instances and accuracy $\mathrm{Acc}(D_{k})$. The micro-averaged accuracy is
+$\mathrm{Acc}_{\text{micro}}=\frac{\sum_{k=1}^{K}|D_{k}|\,\mathrm{Acc}(D_{k})}{\sum_{k=1}^{K}|D_{k}|},$
+which weights datasets by size. The macro-averaged accuracy is
+$\mathrm{Acc}_{\text{macro}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{Acc}(D_{k}),$
+which weights all datasets equally.
+B.3. Insights Learned Through Baseline Experience Learning Methods
+To better understand the limitations of prior experience-learning methods in optimization formulation, we examine all insights extracted by ExpeL when adapted to operations research tasks.
+The learned insights are dominated by generic, high-level recommendations, listed below:
+  • Ensure that decision variables are defined with the correct type (integer, binary, or continuous) based on the problem requirements, especially for inherently discrete quantities.
+  • Double-check that all constraints are correctly implemented and aligned with the problem conditions, including the correct mathematical representation and direction of inequalities.
+  • Verify that the objective function accurately reflects the intended optimization goal.
+  • Always validate the model output against expected results to ensure feasibility and correctness.
+  • Regularly review the problem statement to ensure correct interpretation of variable types and constraints.
+  • Ensure that decision variables have appropriate bounds to reflect realistic limits and prevent infeasible or unbounded solutions.
+  • Incorporate logical constraints to model conditional relationships between decision variables.
+  • Test the model under different scenarios to assess robustness and adaptability.
+  • Re-validate variable types when transitioning between continuous and discrete representations.
+  • Check the model structure for potential infeasibilities or unboundedness before optimization.
+  • Consider edge cases to improve robustness across problem conditions.
+  • Ensure that the objective and constraints scale appropriately with problem size or parameter changes.
+  • Examine the impact of variable bounds on feasibility and correctness.
+These insights emphasize general modeling hygiene rather than concrete modeling or coding practices.
+They do not encode explicit applicability conditions, domain-specific structural patterns (e.g., flow conservation, inventory balance, fixed-charge activation), or solver-level implementation details.
+Consequently, while such insights are broadly reusable, they provide limited actionable guidance for correcting specific formulation or implementation errors in complex optimization problems.
+B.4. Experimental Cost
+All services used in this work rely on third-party API calls to OpenAI’s GPT-4o models.
+We report the token consumption incurred during both the training and evaluation phases of AlphaOPT.
+Here Steps denotes the number of execution rounds required by each module or phase.
+In practice, the iterative Library Diagnosis–Refinement circle in AlphaOPT typically converges within five rounds during training.
+As shown in the table, the Library Refinement phase consumes more prompt tokens, since refining each insight requires incorporating all of its associated tasks as contextual input.
+During the testing phase, the token cost is only related to solution generation and insights resue, resulting in a substantially lower token consumption compared to the training stage.
+Overall, under well-designed prompt engineering and constrained execution steps, the average cost per instance remains around \$ 0.1, and the total experimental cost is kept at a moderate and acceptable level.
+Moreover, owing to the reusability of the library, the relatively high computational cost is incurred only once during the construction of experience library; therefore, the overall system exhibits strong economic scalability in large-scale testing or long-term usage scenarios.
+Table 4. Computation cost and runtime statistics across training and testing stages. (to be updated
+Stage Phase Data Size Prompt Tokens Completion Tokens Cost per instance Runtime Steps
+Training Library Learning 452 $\sim$7.7M $\sim$1.3M $\sim$\$0.07 $\sim$40.5 min 1
+Training Library Diagnosis 452 $\sim$11.5M $\sim$2.3M $\sim$\$0.11 $\sim$48.6 min 3-5
+Training Library Refinement 452 $\sim$32.9M $\sim$1.9M $\sim$\$0.22 $\sim$33.7 min 3-5
+Testing 691 $\sim$13.8M $\sim$1.5M $\sim$\$0.07 $\sim$25.5 min 3
+B.5. Insight Distribution
 To further understand the detailed content of the library insights, we analyze the distributions of insights under the level-2 taxonomy labels, as well as the contribution differences among training datasets (as depicted in Figure 7).
@@
 (c) Insight distribution of Code Implementation track
-Figure 8: Insight distributions across three different tracks, showing the contributions of source tasks in the four training datasets to the generation of library insights.
+Figure 7. Insight distributions across three different tracks, showing the contributions of source tasks in the four training datasets to the generation of library insights.
+The figure shows how insights in the experience library are distributed across the three main tracks and how different source datasets contribute to each track.
+Panel (a) focuses on the Domain Modeling track.
+Each horizontal bar corresponds to a specific domain-level modeling concept, such as capacity and resource balance equations, flow conservation, max-flow objectives, or inventory balance.
+The bars are segmented by dataset, indicating how many insights of each type originate from NLP4LP, Industry OR, MAMO Complex, and NL4OPT tasks.
+This panel highlights that domain modeling insights are concentrated around a small number of common structures, especially resource allocation and network flow patterns, with contributions coming from multiple datasets.
+Panel (b) shows the distribution for the General Formulation track.
+The listed concepts include variable definition issues, constraint formulation errors, unit scaling problems, and objective specification confusions.
+Compared to domain modeling, these insights are more evenly spread across formulation-level pitfalls, such as continuous versus discrete variables, redundant auxiliary variables, and Big-M related issues.
+Contributions again vary across datasets, reflecting differences in how formulation errors arise in different task sources.
+Panel (c) presents the Code Implementation track.
+The insights here relate to solver and API usage, constraint syntax, variable bounds, and data input or validation errors.
+The distribution shows fewer categories but clear concentration on solver syntax and implementation mistakes, with notable contributions from datasets that emphasize executable optimization code.
+Overall, the figure illustrates that different datasets contribute complementary types of insights and that the learned library captures a diverse range of issues spanning domain modeling, mathematical formulation, and code-level implementation.
 In the Domain Modeling track, insights under Resource Allocation – Capacity/Resource Balance Equations account for the largest proportion, with contributions from all four datasets, among which NLP4LP and Industry OR contribute the most.
-Insights tagged with this label are widely applicable to optimization problems that require ensuring resource usage does not exceed available capacity, demonstrating how to establish constraints that maintain balance between resource consumption and availability.
+Insights tagged with this label are widely applicable to optimization problems that require ensuring that resource usage does not exceed available capacity.
+This demonstrates how to establish constraints that maintain balance between resource consumption and availability.
 In the General Formulation track, Variable Definition – Explicit Bounds, Redundant Auxiliary Variable, and Continuous vs.
-Discrete Confusion are the most frequent, with diverse sources, indicating that these are common issues across multiple datasets.
-This reflects that foundational concepts in variable definition are the most error-prone in optimization modeling—particularly in balancing variable types, value ranges, and modeling simplifications—where LLMs tend to produce redundant formulations or type misuse due to overlooking structural or physical consistency.
-In the Code Implementation track, the number of insights is the smallest, with Solver & API Syntax – Nonlinear Constraint Handling and Data I/O & Validation – KeyError & Index Mismatch accounting for the highest proportions, mainly contributed by the MAMO (Complex LP) and NLP4LP datasets.
-These two types of insights reveal critical vulnerabilities in the implementation stage—solver syntax compatibility and data accessibility—representing the dual pillars required for bridging mathematical modeling and executable code.
+Discrete Confusion are the most frequent, with diverse sources, which indicates that these are common issues across multiple datasets.
+This reflects the fact that foundational concepts in variable definition are the most error-prone in optimization modeling—particularly in balancing variable types, value ranges, and modeling simplifications—while LLMs tend to produce redundant formulations or type misuse due to overlooking structural or physical consistency.
+In the Code Implementation track, the number of insights is the smallest, with Solver & API Syntax – Nonlinear Constraint Handling and Data I/O & Validation – KeyError & Index Mismatch accounting for the highest proportions, which are mainly contributed by the MAMO (Complex LP) and NLP4LP datasets.
+These two types of insights reveal critical vulnerabilities in the implementation stage—solver syntax compatibility and data accessibility—and represent the dual pillars required for bridging mathematical modeling and executable code.
 From the perspective of dataset contribution, Industry OR is dominant in Domain Modeling; MAMO (Complex LP) and Industry OR lead in General Formulation; and both MAMO (Complex LP) and NLP4LP contribute the most to Code Implementation.
 NL4OPT has relatively lower overall participation but focuses on formulation- and solver-related details.
-Considering dataset size, although the Industry OR and MAMO (Complex LP) datasets used for library learning are only about half the size of the other datasets, they still contribute a large number of insights, indicating that these datasets contain denser structural modeling challenges and more diverse error patterns, enabling the LLM to accumulate more experiential knowledge across multiple dimensions.
-D.3 Examples of Library Refinement
-We selected five representative refined insights from the library and explained how their applicability conditions were adjusted based on the associated tasks.
-Figure 9: The examples of insight conditions before and after refinement, with green text indicating the original applicability conditions, and blue and red marking the newly added applicable and non-applicable conditions after refinement.
-As shown in the Figure 9, Insight ID 10 targets single–scalar-capacity selection; the refinement broadens coverage by adding equivalent formulations such as minimizing a single resource (time/space), while explicitly excluding multi-capacity settings.
-Insight ID 43 captures flow/conservation and production balancing across nodes; refinement adds lexical anchors (e.g., capacity constraints, balancing production) and excludes statements dominated by fixed proportions, quality constraints, or predefined allocations, gating suppresses spurious retrievals where the structure is not true flow/conservation.
-Insight ID 100 supports adding tight explicit bounds derived from natural limits; refinement strengthens positive cues (e.g., maximum capacity, resource limits) and excludes cases that are purely discrete or already inherently bounded.
-Insight ID 133 focuses on max-flow between a single source and a single sink; refinement tightens the structural requirement and rules out source-to-many distribution tasks.
-Insight ID 228 covers translating at least/at most expressed as proportions into inequalities; refinement broadens to general percentage phrasing, clarifies the mapping, and excludes exact quantities, logical relations, or fixed ratios that are not proportion limits.
-Across the five cases, four strategies proposed by LLM agent recur: (i) generalize with equivalent phrasings (e.g., minimize a single resource); (ii) lexical anchoring with positive keywords to raise recall where appropriate; (iii) explicit exclusion lists to reducing misalignment with tasks; (iv) structural qualifiers (single-scalar capacity; single-source–single-sink) to prevent misuse.
-A refinement is considered effective when the reduction in negative and unretrieved tasks outweighs the decrease in positive tasks.
-As shown in the Figure 10, these semantic refinements reduce both negative and unretrieved cases while preserving as many positive cases as possible.
-Figure 10: Changes in the number of positive, negative, and unretrieved tasks associated with each insight before and after applicability condition refinement.
-A decrease in negative and unretrieved tasks indicates that, the insight no longer mismatches unrelated tasks (negative) and successfully matches previously applicable but missed tasks (unretrieved) after refinement.
-D.4 Specification of Library Taxonomy Labels
-The following Table  6 lists all library taxonomy labels and their corresponding conditions, which specify the applicability criteria of each level-2 label and clarify its precise meaning.
-According to the library taxonomy generation mechanism, each label condition is created by the LLM when the label is first introduced, and subsequent generated insights under the same label inherit that initial condition.
-Table 6: Full Specification of the Library Taxonomy
+Considering dataset size, although the Industry OR and MAMO (Complex LP) datasets used for library learning are only about half the size of the other datasets, they still contribute a large number of insights, which indicates that these datasets contain denser structural modeling challenges and more diverse error patterns; this enables the LLM to accumulate more experiential knowledge across multiple dimensions.
+B.6. Success and Failure Case Study
+To demonstrate the effectiveness of the retrieved library insights, we perform case studies across all evaluation datasets (including the test splits of NLP4OPT, NLP4LP, IndustryOR, and MAMO-complex, as well as the out-of-distribution sets LogiOR and Optibench) in which we compare the performance of solving optimization tasks with and without library retrieval enabled, while excluding tasks that succeed in both settings from the evaluation.
+As shown in Figure 8, approximately half of the retrieved insights successfully contributed to solving new optimization tasks, with only a small portion introducing new errors (Failure).
+This demonstrates the effectiveness of our method for automating knowledge transfer for optimization modeling.
+Moreover, insights under different level-1 taxonomy labels exhibit distinct performance patterns across modeling dimensions.
+(a) Overall effectiveness of retrieved library insights
+(b) Effectiveness distribution across Level-1 insight categories
+Figure 8.
+We evaluate the effectiveness of retrieved insights by classifying their outcomes as success, failure, or invalid.
+An insight is considered successful if it matches a new task and prevents an error that would otherwise occur, failed if it mismatches a task and introduces a new error, and invalid if it matches correctly but fails to help the LLM fix the original mistake.
+(a) The overall distribution of insight outcomes.
+(b) The proportions of successful, failed, and invalid insights under level-1 taxonomy labels.
+The figure evaluates how effective the retrieved library insights are when applied to new tasks.
+Panel (a) shows the overall outcome distribution for all retrieved insights.
+Each insight is classified as a success, failure, or invalid.
+Successful cases are those where the insight matches the task and prevents an error that would otherwise occur.
+Failed cases are those where the insight mismatches the task and introduces a new error.
+Invalid cases are those where the insight is applicable but does not help correct the original mistake.
+The pie chart reports both counts and proportions, showing that nearly half of the retrieved insights lead to successful corrections, while the rest are split between invalid and failed outcomes.
+Panel (b) breaks down these outcomes by level-1 insight categories.
+Each horizontal bar corresponds to a category such as domain modeling, general formulation, or code implementation, and is further divided into success, failure, and invalid portions.
+This view highlights differences in effectiveness across categories.
+For example, variable definition and solver or API syntax insights account for a large number of successful cases, while some domain modeling categories show higher proportions of invalid outcomes.
+Overall, the figure shows that retrieved insights are often helpful, with some variation between categories.
+1) Insights with high success rates
+These insights are concentrated in Code Implementation – Solver & API Syntax and General Formulation – Variable Definition.
+These insights typically target code implementation or fundamental modeling errors with clear structures, which enable the LLM to follow their guidance stably and correctly.
+For instance, insights tagged with the Strict Inequality label under Solver & API Syntax highlight the fact that solvers (e.g., Gurobi) do not support strict inequalities and should instead be reformulated as non-strict forms ($\leq-1$ or $\geq+1$); insights tagged with the Explicit Bounds label under Variable Definition emphasize that decision variables should be assigned explicit upper and lower bounds to ensure feasibility and improve solver efficiency.
+2) Insights with high failure rates
+These insights are mainly found in Domain Modeling – Facility Location.
+They often involve structural constraints and logical triggers that are easily misinterpreted or overgeneralized by the LLM.
+For example, insights tagged with the Fixed Charge (Big-M Linking) label exhibit a failure rate of 70%.
+Although the principle of using Big-M constraints to model on/off logic is correct, its blind application in problems without conditional activation can lead to redundant or overlapping constraints and unnecessary feasible-region reduction, which ultimately causes solver failure.
+Also, while insights under Variable Definition generally achieve high success rates, some insights tagged with Explicit Bounds label sometimes are overly rigid, which leads the LLM to impose unnecessary upper bounds and thereby restricts the feasible space and produces suboptimal solutions.
+3) Insights with high invalid rates
+These insights are mainly under Domain Modeling – Resource Allocation and Domain Modeling – Product Planning.
+Although the LLM successfully retrieves the correct insights and identifies the corresponding problem types, it often fails to translate them into executable formulas or solver-level implementations.
+For instance, under Solver & API Syntax, Nonlinear Constraint Handling advises linearization or the introduction of auxiliary variables for nonlinear objectives or constraints (e.g., ratios or divisions), yet the LLM frequently fails to fully execute these transformations (by neglecting auxiliary variables or mis-rewriting proportional constraints), which results in insights being recognized but not operationalized.
+Also, certain tasks remained unsolved regardless of whether library retrieval was enabled, because their failure stemmed from factors beyond the current learned library’s knowledge scope.
+In the out-of-distribution dataset LogiOR, tasks involved multi-level spatiotemporal logic and interacting constraints (e.g., capacity, timing, and flow balance) in problems such as routing, scheduling, and network flow.
+These challenges extend beyond the scope of the existing library, which primarily focuses on static, linear formulations.
+Although related taxonomy labels such as Resource Allocation and Nonlinear Constraint Handling exist, their granularity and depth are insufficient for modeling such complex logic.
+This demonstrates that the current system, while semantically generalizable, still lacks robust cross-structural transfer and context adaptation capabilities.
+Overall, while successful insights constitute the majority, the results reveal several directions for improvement.
+First, the relatively high failure rates indicate that, despite the inclusion of condition refinement, retrieval precision can still be improved through enhanced semantic disambiguation and structural filtering.
+Second, insights with high invalid proportions suggest the need for clearer explanations and better-designed examples to improve pedagogical clarity and execution effectiveness.
+Finally, for out-of-distribution tasks, future efforts should focus on strengthening the LLM’s ability to adapt and generalize retrieved insights to unseen, complex optimization scenarios.
+Moreover, expanding OR datasets based on LLM error typologies can further enhance experiential learning efficiency and generalization at comparable problem scales.
+B.7. Specification of Library Taxonomy Labels
+The following Table  5 lists all library taxonomy labels and their corresponding conditions, which specify the applicability criteria of each level-2 label and clarify its precise meaning.
+According to the library taxonomy generation mechanism, each label condition is created by the LLM when the label is first introduced, and subsequent insights generated under the same label inherit that initial condition.
+Table 5. Full Specification of the Library Taxonomy
 Taxonomy Label Condition
@@
 Production Planning
-Inventory Balance Equations Applies when the problem description involves materials or products that carry over between periods and must satisfy stock-flow balance.
+Inventory Balance Equations Applies when the problem description involves materials or products that carry over between periods and must satisfy the stock-flow balance.
 Technical Ratio Constraints Applies when the problem description specifies minimum/maximum production ratios or recipe proportions between products or stages.
@@
 Missing Data Defaults Applies when the mathematical model contains optional parameters whose values may be absent in the dataset and require default assignments to preserve model validity.
-Appendix E Prompts For LLM Modules
+Appendix C Prompts For LLM Modules
 Apply self-exploration on finding gold-standard program You are an expert in Industrial Engineering and Operations Research.